### v2

# Magnetic Polarisability of Octet Baryons via Lattice QCD

Abstract
Drawing on recent advances in lattice-QCD background-field techniques, the magnetic polarisability of octet baryons is calculated from the first principles of QCD. The results are presented in the context of new constituent quark-model calculations providing a framework for understanding the lattice results and a direct comparison with simulation results at unphysical quark masses. Using smeared quark sources, low-lying Laplacian eigenmode projection and final-state Landau mode projection, considerable attention is devoted to ensuring single-state isolation in the lattice correlation functions. We also introduce new weighting methods to reduce the sensitivity to correlation-function fits, averaging over many fits based on merit drawn from the full correlated $\chi^{2}$ of the fits. The techniques are implemented on the $32^{3}\times 64$, $2+1$-flavour dynamical-fermion lattices provided by the PACS-CS collaboration following the introduction of uniform magnetic fields quantised to the lowest nontrivial values available. After some fine tuning of the constituent quark model parameters, we find the model captures the patterns observed in the lattice QCD results very well, providing important insights into the physics underpinning the magnetic polarisabilities. Finally, comparison with the most recent results from experiment proceeds through an effective field theory formalism which incorporates estimates of finite-volume corrections and small electro-quenching corrections as the results are brought to the physical point. We find excellent agreement with experiment where available, including the proton and neutron polarisabilities.
pacs:
13.40.-f, 12.38.Gc, 12.39.Jh
I Introduction
The magnetic polarisability is the second order response of an electrically charged composite particle to an external magnetic field. It encapsulates the manner in which the internal structure of the particle is changed by the field. From a perturbative point of view, the second-order process induces virtual electromagnetic transitions to nearby excitations in the hadron spectrum and probes the distribution of quarks within the hadron. In a nonperturbative sense, these virtual transitions act to change the structure of the hadron and thus the energy of the particle as it resides in the magnetic field.
Our focus is on lattice QCD calculations of the magnetic polarisability of octet baryons. We draw on recent advances in lattice-QCD background-field techniques, established in an analysis of nucleon polarisabilities Bignell et al. (2020). In particular, smeared quark sources are utilised to capture the QCD aspects of the hadron structure. Then low-lying QCD+QED SU(3)$\times$×U(1) Laplacian eigenmode projection is considered at the quark level to capture the low-energy response of individual electrically-charged quarks to QCD and the external magnetic field. Electrically charged baryons will be in a superposition of Landau-level states and therefore final-state colour-singlet U(1) Landau-mode projection is used at the sink to ensure the baryon is in the lowest-lying Landau level Bignell et al. (2018, 2020); Tiburzi and Vayl (2013). We extend the previous analysis by devoting considerable attention to ensuring single-state isolation in the lattice correlation functions. We further introduce new weighting methods to reduce the sensitivity to correlation-function fits, averaging over many fits based on merit determined by the full covariance-matrix based $\chi^{2}$ of the fits.
To help in understanding the pattern of octet-baryon magnetic polarisabilities observed, we consider the simple nonrelativistic constituent quark model and extend calculations of the nucleon polarisabilities Capstick and Keister (1992); Lipkin and Moinester (1992); Bhaduri and Warke (1988) to include the hyperons of the baryon octet. As we will demonstrate, chiral nonanalytic behaviour is subtle at the quark masses we consider on the lattice. Therefore we anticipate this simple model will be of utility in understanding the competing effects that generate some complexity in the pattern of magnetic polarisabilities observed in lattice QCD. Qualitatively, it explains the large difference between negatively charged baryons and the remainder of the octet. After some fine tuning of the constituent quark model parameters, we find the model captures the patterns observed in the lattice QCD results very well, providing important insights into the physics underpinning the magnetic polarisabilities.
Our approach to the magnetic polarisability in lattice QCD centres on the uniform background magnetic field method. Historically, this approach has been the method of choice for determining baryon magnetic polarisabilities Zhou et al. (2003); Lee et al. (2006); Primer et al. (2014); Chang et al. (2015); Bignell et al. (2018, 2020). As a leading effect in the expansion of the baryon energy in terms of the magnetic field strength, the background field must be weak to ensure higher ${\cal O}(B^{3})$ terms are small. Thus, subtle shifts in the energy of the baryon induced by the magnetic field need to be related to the polarisability. To reveal the effects of the magnetic field, correlated ratios are constructed to enable QCD based fluctuations to cancel.
Our techniques are implemented on the $32^{3}\times 64$, $2+1$-flavour dynamical-fermion lattices provided by the PACS-CS collaboration. QCD correlations are maintained by applying the external magnetic field to these existing gauge field configurations. A consequence of this is that our gauge fields are electro-quenched in that the sea-quarks are blind to the magnetic field. At the quark masses considered herein, effective field theory indicates the approximation is reasonable, and we use effective field theory to estimate the small electro-quenching corrections Deshmukh and Tiburzi (2018); Hall et al. (2014); Bignell et al. (2020). We also draw on effective field theory to estimate the small finite-volume corrections to our lattice QCD calculations Hall et al. (2014); Bignell et al. (2020).
In constructing the required correlated ratios, several two-point correlation functions are combined. These include the zero-field correlator and spin-field aligned and anti-aligned baryon correlators at finite magnetic-field strength. These all display different behaviour as a function of Euclidean time and their combination can hide excited-state contamination effects. To address this, we examine the Euclidean-time behaviour of each of these underlying correlators to ensure that the effective energy of each two-point correlator has reached single-state isolation through Euclidean time evolution. We find that in many cases plateau-like behaviour in the correlated ratios precedes the single-state isolation of the individual underlying correlation functions. Thus the results presented herein provide an important step forward in suppressing contamination from excited states.
We also carefully examine the dependence of the results on the selection of the fit window where a significant dependence can arise in some cases. To address this we introduce a weighting method to reduce the sensitivity to fit-window selection. Following the ideas presented in Ref. Beane et al. (2021), we average over many fit windows with their weight based on merit calculated with the full covariance-matrix based $\chi^{2}$ of the fits.
We find the processes described above bring a new level of rigour to the polarisability results, producing the highest precision lattice QCD calculations of baryon magnetic polarisabilities to date.
The presence of a magnetic field and the difference in the $u$ and $d$ quark charges breaks the isospin symmetry of equal mass quarks. The $\Lambda$ and $\Sigma^{0}$ baryons will mix in the magnetic field complicating the interpretation of the mass shifts in the background field approach considered herein. Correlation matrix techniques can be introduced to address the state isolation issue. However, together these issues add additional layers of complexity that is not required for the other six members of the baryon octet. These are the so-called outer members of the octet, composed with a doubly-represented quark flavour and a singly-represented flavour. It is these baryons that hold the focus of the present investigation.
Recently four-point function methods have been considered in the calculation of the magnetic polarisability Wilcox and Lee (2021); Lee et al. (2023). This time the baryons are probed perturbatively by two electromagnetic-current insertions. As such, the baryon states are eigenstates of QCD alone and the $\Lambda$ and $\Sigma^{0}$ baryons do not mix. Thus, this alternative approach presents a significant advantage to understanding the magnetic polarisabilities of the $\Lambda$ and $\Sigma^{0}$ baryons and we anticipate future calculations for these baryons.
The presentation of this work is as follows. In Sec. II we derive the generalised expression for the magnetic polarisability of an octet baryon in a simple constituent quark model. Highlighting the competing effects of magnetic transitions and the distribution of charge within the baryon, we derive predictions for the magnetic polarisabilities of the outer octet baryons and use these to provide context for the results of the lattice QCD calculations. In Sec. III we describe our implementation of the background field for the lattice QCD calculation, emphasising the co-location of the field and baryon correlator origin. In Sec. IV we discuss the extraction of the magnetic polarisability from two-point baryon correlation functions and ratios. In Sec. V we discuss relevant details of the lattice simulation including the ensembles used, the quark level SU(3)$\times$×U(1) projection at the quark-propagator sink, and the U(1) hadronic projection which allows for proper handling of Landau-level physics. In Sec. VI we discuss our advances in the correlation-function fitting methods, including the examination of the excited state contamination in the underlying correlation functions and the weighted averaging of fit windows. In Sec. VII we present the results of the lattice calculation and utilise them to improve the quark model such that it can capture the essential physics governing baryon magnetic polarisabilities. In Sec. VIII we use chiral effective field theory to incorporate electro-quenching corrections and estimate finite-volume corrections before extrapolating the lattice results to the physical regime. In Sec. IX we compare our findings with experimental measurements and previous determinations of octet-baryon magnetic polarisabilities. In Sec. X we briefly examine the proton-neutron magnetic polarisability difference. In Sec. XI we summarise our findings.
II Constituent Quark Model
The constituent quark model is renowned for its capacity to provide a simple explanation for the pattern of baryon magnetic moments observed in experiment. Founded on SU(6) spin-flavour symmetry, symmetric ground-state wave functions with zero angular momentum, and anti-symmetry via SU(3) colour-singlet states, the model captures the essence of the physics with subtle corrections arising from the environment sensitivity of the constituent quark moments Leinweber et al. (1991); Boinepalli et al. (2006). We consider the constituent quark model predictions for octet-baryon magnetic polarisabilities in a similar spirit, hoping to capture the predominant features in a simple model and providing a framework for the description of more subtle corrections.
II.1 Magnetic polarisabilities from the quark model
Bhaduri et al. Bhaduri and Warke (1988) present an expression for the magnetic polarisability of the proton in the constituent quark model. Here we review the derivation of this expression and generalise it for an arbitrary octet baryon. In doing so, we aim to provide insight into the salient features of QCD that give rise to the magnetic polarisabilities of ground-state baryons, at quark masses where pion dressings of the baryons are of secondary importance Hall et al. (2014); Bignell et al. (2020).
A quark model expression for the magnetic polarisability may be obtained using Rayleigh-Schrödinger perturbation theory, reviewed in Ref. Sakurai and Commins (1995) for example. We consider an unperturbed Hamiltonian $\mathcal{H}_{0}$, and label the energy eigenstates of this Hamiltonian |$\ket{n_{i}}$
$\mathcal{H}_{0}\ket{n_{i}}=E_{i}\ket{n_{i}},\quad i=1,2,3,\ldots\,.$ ((1))
This Hamiltonian is perturbed with an interacting Hamiltonian $\mathcal{H}_{\rm int}$. In our case this will be a Hamiltonian relevant to the inclusion of a magnetic field. Substituting this perturbation into the Schrödinger equation we define
$(\mathcal{H}_{0}+\lambda\mathcal{H}_{\rm int})\ket{k_{i}}=E^{\prime}_{i}\ket{k_{i}},\quad i=1,2,3,\ldots\,.$ ((2))
where |$\ket{k_{i}}$ are the energy eigenstates of the full Hamiltonian and $\lambda$ controls the strength of the perturbation. At second order in $\lambda$, the perturbed energy of the $i$th energy level is
$E^{\prime}_{i}(\lambda)$ $=$ $E_{i}+\lambda\bra{n_{i}}\mathcal{H}_{\rm int}\ket{n_{i}}$ ((3))
[…]
and the shift in energy
$\Delta E=E^{\prime}_{i}(\lambda)-E_{i}\,,$ ((4))
includes terms both linear and quadratic in $\lambda$. At this point we set $\lambda=1$ and draw on the electric charge, $e$, in $\mathcal{H}_{\rm int}$ to ensure the interaction is perturbative.
The magnetic polarisability is associated with terms quadratic in the magnetic field strength and thus $e^{2}$ and we’ll see both correction terms of Eq. (3) provide a contribution. Focusing on terms proportional to $e^{2}$, we define the magnetic polarisability, $\beta$, via
[…]
where $\Delta E_{2}$ contains the terms second order in $e$ from Eq. (3).
In deriving the interacting Hamiltonian, we first consider the case of a free particle, where we can write the Hamiltonian in terms of the momentum operator and the mass
[…]
Minimal substitution provides Anderson (1971)
$\vec{p}\rightarrow\vec{p}+q\,e\,\vec{A}\,,$ ((7))
where $q\,e$ is the particle charge and $\vec{A}$ is the electromagnetic vector potential. Making this substitution, we obtain
[…]
where we note that $\vec{p}$ and $\vec{A}$ commute when the background field is uniform. For $\vec{B}$ in the $z$-direction, $\vec{B}=B\,\hat{k}$, the following relation equitably distributes the field strength among the components of $\vec{A}$
$\vec{A}$ $=\frac{1}{2}\,\vec{B}\crossproduct\vec{r}\,,$ ((9))
$=\frac{1}{2}(-B\,y,B\,x,0)\,,$ ((10))
This relation allows us to rewrite the dot product in terms of angular momentum
$\vec{p}\cdot\vec{A}$ $=\frac{1}{2}\,\vec{p}\cdot\left(\vec{B}\times\vec{r}\right)\,,$ ((11))
$=\frac{1}{2}\,\vec{L}\cdot\vec{B}\,.$ ((12))
Here we are purely interested in ground state baryons. As such, the quarks are in a relative S-wave state and have zero angular momentum causing that term to vanish, leaving
[…]
We must also account for the particle’s spin. Inclusion of a vector potential in the Dirac equation for a charged spinor produces a term Schwartz (2014)
$\frac{q\,e}{2m}\,\vec{B}\cdot\hat{\sigma}\,,$ ((14))
where $\hat{\sigma}$ is the Pauli spin operator. As such our Hamiltonian becomes
[…]
where the magnetic moment operator is given by
$\hat{\mu}=\frac{q\,e}{2m}\,\hat{\sigma}\,.$ ((16))
The sign of the magnetic moment term comes from the fact that an addition of a magnetic field will cause the particle’s spin to align opposite to the field, thereby decreasing its energy via the dot product.
Drawing on Eq. (10),
[…]
and
[…]
where we now see that the $e^{2}$ term contains the position operators. Thus the magnetic polarisability contains contributions probing the distribution of quarks within the hadron.
Recalling $\vec{B}=B\,\hat{k}$, we reduce the dot product to simply the magnetic moment operator acting purely in the $z$-direction. Further, as we are working within the framework of the quark model we include a sum over the quarks in our baryon. As such, we write our interaction Hamiltonian as
$\mathcal{H}_{\rm int}=\hat{\mu}_{z}\,B+\sum_{f=1}^{3}\frac{q_{f}^{2}\,e^{2}}{8m_{f}}(\hat{x}^{2}+\hat{y}^{2})B^{2}\,,$ ((19))
where our magnetic moment operator is now
$\hat{\mu}_{z}=\sum_{f=1}^{3}\frac{q_{f}\,e}{2m_{f}}\,\hat{\sigma}_{z}\,.$ ((20))
This interaction Hamiltonian is then substituted into Eq. (3). Such a substitution requires the determination of two matrix elements. Firstly
|$\bra{\mathcal{B}}\mathcal{H}_{\rm int}\ket{\mathcal{B}}$ $=\mu_{\mathcal{B}}\,B+\sum_{f=1}^{3}\frac{q_{f}^{2}\,e^{2}}{8m_{f}}\left(\expectationvalue{x^{2}}_{f}+\expectationvalue{y^{2}}_{f}\right)\,,$
$=\left(\frac{4}{3}\mu_{D}-\frac{1}{3}\mu_{S}\right)B+\sum_{f=1}^{3}\frac{q_{f}^{2}\,e^{2}}{12m_{f}}\expectationvalue{r^{2}}_{f}\,,$ ((21))
where $\mu_{\mathcal{B}}$ is the magnetic moment of the baryon, and $\mu_{D}$ and $\mu_{S}$ are the intrinsic magnetic moments of the doubly-represented and singly-represented quark flavours respectively. In the weak field limit, spherical symmetry of the $S$-wave states provides $\expectationvalue{x^{2}}=\expectationvalue{y^{2}}=\frac{1}{3}\expectationvalue{r^{2}}$.
The second matrix element of interest in Eq. (3) contains an exited state of the octet baryon in the ket. The sum is estimated by considering dominance of the first nearby excitation, the corresponding decuplet baryon state. This matrix element generates terms proportional to $e^{2}$ via
[…]
This contribution is the transition magnetic moment for an octet baryon to its complementary decuplet baryon. For a baryon on the outer ring of the octet, the quark model provides
|$\bra{\mathcal{B}}\hat{\mu}_{f}\ket{\mathcal{B}^{*}}=\mu_{\mathcal{B}\mathcal{B}^{*}}=\frac{2\sqrt{2}}{3}\left(\mu_{D}-\mu_{S}\right)\,,$ ((23))
where $D$ and $S$ once again label the constituent quark flavours for the doubly- and singly-represented sectors; for example for the proton, $\mu_{D}-\mu_{S}=\mu_{u}-\mu_{d}$.
Substituting these matrix elements into Eq. (3), the energy shift quadratic in $e^{2}$ and thus $B^{2}$ is
[…]
$=$ 12 $-\frac{1}{2}\,4\pi\,\beta\,B^{2}$
where the definition of Eq. (5) has been used. Thus, the constituent quark model prediction for the magnetic polarisability of octet baryon $\mathcal{B}$ composed of three quark flavours denoted by $f$ is
[…]
$\equiv\beta_{1}-\beta_{2}\,,$ ((25))
where $\alpha={e^{2}}/{4\pi}$ is the fine structure constant. The magnetic polarisability has its origin in two competing terms which we have defined as $\beta_{1}$, a magnetic transition term, and $\beta_{2}$ probing the distribution of quarks within the baryon. The two terms and their opposing signs in this expression highlight the complex nature of the baryon polarisability. This is the constituent quark model description for the magnetic polarisability of an octet baryon with two quarks of one flavour and another quark with a different flavour.
Using Eqs. (21) and (23) it can be shown that $\bra{p}\hat{\mu}_{z}\ket{\Delta}=\frac{2\sqrt{2}}{3}\mu_{p}$ for constituent quark masses $m_{u}=m_{d}$. Upon substitution into Eq. II.1, one obtains the original result of Bhaduri et al. Bhaduri and Warke (1988)
$\beta_{p}=\frac{16}{9}\frac{1}{4\pi}\frac{\mu_{p}^{2}}{m_{\Delta}-m_{p}}-\frac{\alpha}{6\,m_{l}}\left(2\,q_{u}^{2}\expectationvalue{r^{2}}_{u}+q_{d}^{2}\expectationvalue{r^{2}}_{d}\right),$ ((26))
where $m_{l}=m_{u}=m_{d}$.
To compare the predictions of the quark model to the results of our lattice QCD calculations, we require a relationship between the baryon masses calculated in lattice QCD and the associated constituent quark mass to be used in the model. We turn our attention to this in the following section.
II.2 Implementing the quark model
To present the predictions of the quark model for the magnetic polarisability as a function of quark mass, we require a relationship between the lattice baryon masses and the constituent quark masses. This relationship has been long established for the baryon octet at the experimental point and therefore we need only extend this relationship away from the physical point. In addition, the octet-decuplet mass splitting in Eq. II.1 demands knowledge of both octet and decuplet baryons. Similarly, $\expectationvalue{r^{2}}_{f}$ describing the distribution of quark flavours within the octet baryons is also required. We seek smooth, interpolated expressions for each of these quantities.
Our lattice calculations are performed on the PACS-CS gauge-field ensembles Aoki et al. (2009) and therefore we draw on their published baryon masses in the Sommer scheme. Ref. Walker-Loud (2014) demonstrated that the nucleon masses observed in finite-volume lattice QCD simulations display a ruler-style linear behaviour when plotted as a function of $m_{\pi}$. This provides a simple characterisation of lattice baryon masses and in the spirit of the simple quark model, we take this approach to interpolate the PACS-CS baryon masses for the octet and decuplet.
To estimate the constituent quark masses away from the physical point, we utilise a simple linear model
$m^{0}_{\rm oct}+\alpha_{\rm oct}\,(3\,m_{l})$ $=m_{N},$
$m^{0}_{\rm oct}+\alpha_{\rm oct}\,(m_{l}+2\,m_{h})$ $=m_{\Xi},$ ((27))
which leverages the maximum difference in strange quarks in the baryon octet. Here $m^{0}_{\rm oct}$ and $\alpha_{\rm oct}$ are fit parameters, and $m_{l}$ and $m_{h}$ are the light and heavy constituent quark masses already determined in the simple constituent quark model summarised in Ref. Workman et al. (2022). The parameter $m^{0}_{\rm oct}$ allows some of the nucleon mass to have its origin in the confining potential and $\alpha_{\rm oct}$ allows for spin-dependent effects to change the slope from 1. Using experimental baryon masses and the constituent quark masses $m_{l}=338\,$MeV and $m_{h}=510\,$MeV from the PDG Workman et al. (2022), we solve for the fit parameters $m^{0}_{\rm oct}$ and $\alpha_{\rm oct}$.
For decuplet baryons, we repeat the analysis with two new fit parameters
$m^{0}_{\rm dec}+\alpha_{\rm dec}\,(3\,m_{l})$ $=m_{\Delta},$
$m^{0}_{\rm dec}+\alpha_{\rm dec}\,(3\,m_{h})$ $=m_{\Omega}.$ ((28))
The fit parameters corresponding to the two models are shown in Tab. 1. We see that in the octet case the slope parameter $\alpha_{\rm oct}=1.09$ is larger than unity. This is in accord with anticipated hyperfine effects. Because the strength of the attractive hyperfine interaction is inversely proportional to the product of the constituent quark masses, the attraction diminishes with increasing quark mass. Thus the baryon mass grows at a rate exceeding the rate of the quark mass increase.
On the other hand, the hyperfine interaction is repulsive in decuplet baryons. This time the repulsion diminishes with increasing mass such that the decuplet slope is expected to be less than unity. The value of $\alpha_{\rm dec}=0.85$ is in accord with these expectations.
Table 1: Fit parameters for the simple models of Eq. II.2 and Eq. II.2.
Baryon $\alpha$ $m_{0}$ (MeV)
Octet 1.09 $-169$
[…]
We use the octet fit parameters and the ruler-style interpolated $N$ and $\Xi$ masses to obtain the constituent quark masses as a function of $m_{\pi}$. These masses are given in Tab. 2. We see the expected approach to the physical constituent quark mass 338 MeV Workman and Others (2022) while the strange quark mass remains stable as expected.
Table 2: Constituent quark masses for the PACS-CS gauge field ensembles Aoki et al. (2009). Constituent masses are obtained from the model Eq. II.2 with $(\alpha_{\rm oct},m^{0}_{\rm oct})=$($1.09,-169\,$MeV). All masses are in MeV.
kappa $m_{\pi}^{\rm PACS-CS}$ $m_{l}$ $m_{h}$
0.13700 701 480 528
0.13727 570 447 528
0.13754 411 402 529
0.13770 296 365 529
0.13781 156 328 529
With the constituent quark masses $m_{l}$ and $m_{h}$ determined as a function of $m_{\pi}$, one can use the form of Eqs. II.2 to obtain the $\Sigma$-baryon masses by counting two light quarks and one strange quark
$m_{\Sigma}=m^{0}_{\rm oct}+\alpha_{\rm oct}\,(2\,m_{l}+m_{h})\,.$ ((29))
A similar approach is used to get the interpolated decuplet baryon masses. Drawing on Eqs. II.2 and modifying them to count the number of light and strange quarks, the decuplet baryon masses are obtained within the constituent quark model framework in a consistent manner.
Referring back to Eq. II.1 for the magnetic polarisability, the intrinsic quark magnetic moments contained within the octet-decuplet baryon magnetic transition matrix element are simply
$\mu_{f}=\frac{q_{f}\,e}{2m_{f}}.$ ((30))
Quark distribution radii on the PACS-CS ensembles have been determined by Stokes et al. Stokes et al. (2020) where they examined the proton and neutron. We interpolate the proton and neutron squared radii as linear in $\log(m_{\pi})$ and this provides an excellent description of those lattice results. Eq. II.1 requires quark distribution radii for single quark flavours of unit charge, as the charge factors appear elsewhere in the expression. We define radii for the doubly, $D$, and singly, $S$, represented quark sectors for single quark flavours of unit charge by
$\expectationvalue{r^{2}}_{p}$ $=\phantom{-}2\,\frac{2}{3}\,\expectationvalue{r^{2}}_{D}-\frac{1}{3}\expectationvalue{r^{2}}_{S}\,,$ ((31))
$\expectationvalue{r^{2}}_{n}$ $=-2\,\frac{1}{3}\,\expectationvalue{r^{2}}_{D}+\frac{2}{3}\expectationvalue{r^{2}}_{S}\,,$ ((32))
where charge and quark-counting factors are explicit. Then
$\expectationvalue{r^{2}}_{D}$ $=\frac{1}{2}\left(2\expectationvalue{r^{2}}_{p}-\expectationvalue{r^{2}}_{n}\right),$ ((33))
$\expectationvalue{r^{2}}_{S}$ $=\expectationvalue{r^{2}}_{p}-2\expectationvalue{r^{2}}_{n}.$ ((34))
We approximate the doubly and singly represented strange-quark distribution radii to be that of the light quark at the heaviest PACS-CS quark mass where the light quark mass is similar to that of the strange quark as seen in  Tab. 2.
Figure 1: Quark model predictions for the magnetic polarisabilities of octet baryons are plotted as a function of the squared pion mass as a proxy for the quark mass. Dashed lines represent the pion masses of the PACS-CS ensembles. The legend is ordered to match the vertical ordering at the physical point.
With the octet and decuplet baryon masses, constituent quark masses, magnetic moments, and quark flavour distribution radii parameterised as a function of $m_{\pi}^{2}$, we are now able to present values for the magnetic polarisability as predicted by the constituent quark model as illustrated in Fig. 1. We note that the predicted magnitudes of the proton and neutron magnetic polarisabilities are somewhat larger than expected from experimental measurements. While we will address this issue in the following, the trends of the model are worthy of discussion.
First, we note the quark model predicts qualitatively different values for the negatively charged baryons. Their polarisabilities are predicted to be opposite in sign and much smaller in magnitude. Recalling Eq. II.1 with the transition magnetic moment written explicitly in terms of quark moments using equation Eq. (23)
$\beta$ $=\frac{4}{9\pi}\sum_{\mathcal{B}^{*}}\frac{\left(\mu_{D}-\mu_{S}\right)^{2}}{E_{\mathcal{B}^{*}}-E_{\mathcal{B}}}-\sum_{f=1}^{3}\frac{q_{f}^{2}\,\alpha}{6m_{f}}\expectationvalue{r^{2}}_{f},$
$\equiv\beta_{1}-\beta_{2},$ ((35))
we first highlight the squared-difference between the magnetic moments of each sector
$\left(\mu_{D}-\mu_{S}\right)^{2}=\left(\frac{q_{D}\,e}{2m_{D}}-\frac{q_{S}\,e}{2m_{S}}\right)^{2}.$ ((36))
When the quark charges of the sectors are opposed, such as in the proton, these quark moments sum to produce a large contribution. Conversely, in the case of the negative baryons $\Xi^{-}\,$and $\Sigma^{-}\,$, the $d$ and $s$ quarks both carry a charge of $-1/3$ providing cancellation between the terms. However, the splitting of the $d$ and $s$ quark masses admits a small contribution. It is this cancellation that gives rise to the very small nature of the magnetic polarisabilities of the negative baryons. Further cancellation comes from the opposite sign contribution of $\beta_{1}$ and $\beta_{2}$. It’s clear the $\beta_{2}$ term describing quark distributions dominates the sum.
Conversely, it is the opposition of quark sector electric charges that is the main generator of large magnitudes for the magnetic polarisability. Moreover, this is achieved through the inclusion of a $u$ quark whose charge magnitude is double that of the $d$ or $s$ quarks. When added and squared the contribution of opposing charge pairs, $u-d$ or $u-s$, is an order of magnitude larger than the contribution of a single $d$ or $s$ quark flavour.
It is also insightful to consider the explicit constituent quark mass dependence of both terms. We have
$\beta_{1}\propto\frac{1}{m^{2}_{f}},\quad\beta_{2}\propto\frac{1}{m_{f}}\,.$ ((37))
Given that the quark distribution radius, $\expectationvalue{r^{2}}_{f}$, is also smaller for the strange quark, together these effects suppresses the magnitude of the contributions of the strange quark.
As a result, the polarisability is very sensitive to the up quark. The up quark generates a large $\beta_{1}$, but having two up quarks causes a large opposite contribution from $\beta_{2}$. This explains why the $\Xi^{0}\,$and $n$ have greater magnetic polarisabilities than the $\Sigma^{+}\,$and proton respectively.
Another observation is that the mass splitting between the octet and decuplet baryons in $\beta_{1}$ benefits the hyperon magnetic polarisabilities over the nucleon polarisabilities. In the nucleon, the scalar diquark is composed of two light quarks. The strength of this hyperfine interaction is inversely proportional to the product of the constituent quark masses; hence the hyperfine attraction is strong in the nucleon and the hyperfine repulsion is strong in the $\Delta$. Together, these hyperfine interactions produce a large mass splitting in the magnetic transition term, $\beta_{1}$. In the hyperons considered herein, the scalar diquark is composed of a light $u$ or $d$ quark and a strange quark. This time the magnitude of the hyperfine interaction is weaker and the octet-decuplet mass splitting is smaller. Thus the magnitude of the magnetic transition term, $\beta_{1}$, is enhanced. This effect explains why the magnitude of the $\Xi^{0}\,$polarisability exceeds that of the neutron and would suggest the $\Sigma^{+}\,$exceeds the proton.
The alternate ordering of the magnetic polarisability of the proton and $\Sigma^{+}\,$highlights the complexity of the magnetic polarisability. Here the larger mass of the strange quark causes a reduction in the transition magnetic moment which outweighs that of the octet-decuplet mass splitting and the reduced magnitude of $\beta_{2}$.
The $\Xi^{0}\,$and $\Sigma^{+}\,$both have a $u$-$s$ scalar diquark and therefore $\beta_{1}$ is similar. The difference in the total polarisability is associated with $\beta_{2}$ where the second $u$ quark in $\Sigma^{+}\,$makes a larger negative contribution to the polarisability of the $\Sigma^{+}\,$compared to the second $s$ quark in the $\Xi^{0}\,$whose magnitude is small due to both the charge and the mass of the strange quark.
Finally, we note that the magnetic polarisability of the $\Xi^{0}\,$rises quite steeply as the pion mass approaches the physical point. For all baryons, both $\beta_{1}$ and $\beta_{2}$ increase in magnitude as the physical pion mass is approached. However, for the $\Xi^{0}\,$the increase in $\beta_{1}$ significantly outpaces the increase in $\beta_{2}$ near the physical point. It is a combination of all aspects discussed thus far that make this possible. The magnetic transition term, $\beta_{1}$ benefits from both the presence of a $u$ quark enhancing the numerator, but also the presence of a $u$-$s$ scalar diquark suppressing the mass splitting and thus further enhancing $\beta_{1}$. Moreover, the presence of two $s$ quarks ensures $\beta_{2}$ is a small as possible, through the small electric charge, the suppression factor of the strange quark mass and the reduced distribution radius.
We now progress to the lattice QCD calculation to evaluate the veracity of these predictions.
III Background Field Implementation
The background field method is a well established approach to calculating the magnetic polarisability in a lattice QCD calculation. The discussion here is based on Ref. Primer et al. (2014) founded on Ref. Smit and Vink (1987).
We commence by considering the continuum formulation, where the covariant derivative is modified by the addition of a minimal coupling
$D_{\mu}=\partial_{\mu}+i\,g\,G_{\mu}+i\,q_{f}\,e\,A_{\mu}\,,$ ((38))
where $G_{\mu}$ is the gluonic four-potential with coupling $g$, and $A_{\mu}$ is the electromagnetic four-potential with quark-charge coupling $q_{f}\,e$. On the lattice this addition corresponds to a multiplication of the QCD gauge links by an exponential phase factor
$U_{\mu}(x)\rightarrow U_{\mu}(x)\,e^{i\,a\,q_{f}\,e\,A_{\mu}(x)}\,,$ ((39))
where $a$ is the lattice spacing. To obtain a uniform magnetic field along the $z$-axis we use $\vec{B}=\vec{\nabla}\times\vec{A}$, such that
$B_{z}=\partial_{x}A_{y}-\partial_{y}A_{x}\,.$ ((40))
We exploit the second term with $A_{x}=-B\,y$ to produce a constant magnetic field of magnitude $B$ in the $z$-direction.
The resulting field can be calculated by examining a single plaquette in the $(\mu,\nu)=(x,y)$ plane as
$\Box_{\mu\nu}(x)=\exp\left(i\,q_{f}\,e\,a^{2}\,F_{\mu\nu}(x)\right)\,.$ ((41))
This relation is exact for a constant background field as all higher order terms involve a second or higher order derivative. The plaquette at coordinates $x,y$
$\Box_{\mu\nu}(x,y)$ $=$ $\exp(-i\,a\,q_{f}\,B\,y)\exp(i\,a\,q_{f}\,B(y+a))$ ((42))
$=$ $\exp(i\,a^{2}\,q_{f}\,e\,B)\,,$
giving the desired field away from the $y$-direction boundary. However, on a finite lattice with sites labelled $(0\leq x/a\leq N_{x}-1)$ and $(0\leq y/a\leq N_{y}-1)$, there is a discontinuity at the $y$-boundary due to the periodic boundary conditions used for the QCD fields. To address this, we make use of the first term from Eq. (40), $\partial_{x}A_{y}$ and assign
$A_{y}(x,y)=\begin{cases}0\,,&\mathrm{for\ }y/a<N_{y}-1\,,\\ N_{y}\,B\,x\,,&\mathrm{for\ }y/a=N_{y}-1\,.\end{cases}.$ ((43))
This ensures that the discontinuity at the boundary in the $y$-direction at $y/a=N_{y}-1$ is compensated via the $y$-boundary contribution from $A_{y}$.
Having used both terms available in Eq. (40), the double boundary, $x/a=N_{x}-1$ and $y/a=N_{y}-1$, gives rise to a quantisation condition for the field strength. Here the plaquette takes the value
$\Box_{\mu\nu}(x,y)=\exp\left(i\,a^{2}\,q_{f}\,e\,B\right)\,\exp\left(-i\,a^{2}\,N_{x}\,N_{y}\,q_{f}\,e\,B\right)\,,$ ((44))
which provides the required value when the second exponential $\exp(-i\,a^{2}\,N_{x}\,N_{y}\,q_{f}\,e\,B)=1$. Setting $a^{2}\,N_{x}\,N_{y}\,q_{f}\,e\,B$ to be an integer multiple of $2\pi$ provides the field-strength quantisation
$q_{f}\,e\,B=\frac{2\pi\,n}{N_{x}\,N_{y}\,a^{2}}\,,$ ((45))
where $n$ is an integer specifying the field strength in multiples of the minimum field strength quantum.
Throughout this work, we will specify the field quanta in terms of the charge of the down quark. The quantisation condition becomes
$e\,B=\frac{2\,\pi}{N_{x}\,N_{y}\,a^{2}}\frac{1}{q_{d}}\,k_{d},$ ((46))
such that a field corresponding to integer $k_{d}=1$ is oriented in the negative $\hat{z}$-direction.
While in principle the fermion propagator source can be placed anywhere within the spatial volume, one obtains the optimal signal to noise in the baryon correlation functions when the source is placed at the origin of the electromagnetic potential Primer et al. (2014). Hence as one increases statistics by considering many spatial source positions on a single gauge field, we cycle the selected fermion source to the (0,0,0) position by periodic circular shifts of the gauge field. Then the magnetic field is introduced as a phase on the gauge field links as described above.
IV Lattice QCD Formalism
In the presence of a uniform background magnetic field, the energy of a baryon changes as a function of magnetic field strength Martinelli et al. (1982); Primer et al. (2014)
[…]
Here the mass of the baryon, $m$, is complemented by contributions from the magnetic moment $\vec{\mu}$, the Landau term proportional to $|q_{B}\,e\,B|$ where $q_{B}$ is the charge of the baryon, and the magnetic polarisability $\beta$. For neutral baryons, the Landau term does not contribute, and for charged baryons, we will use a U(1) Landau-mode projection to select $n=0$. This aspect of the calculation is detailed in subsection V.4.
This expression arises as a non-relativistic Taylor expansion of the relativistic energy Itzykson and Zuber (2006). The large mass scale of the octet baryons enables the use of the non-relativistic approximation with systematic errors of ($\order{1\%}$ ) at the field strengths considered in this work. See for example, Chapter 3.6.2 of Ref. Bignell (2020). This approach simplifies the required calculations and improves the signal relevant to the polarisabilities.
The magnetic polarisability may be extracted by isolating the quadratic $B^{2}$ term in Eq. (47). The sign of the magnetic moment term is dependent on the alignment of the baryon’s spin, quantised in the $z$-direction, and the magnetic field. Hence it may be eliminated by summing energies associated with different spin orientations
$E_{\uparrow\uparrow}(B)+E_{\uparrow\downarrow}(B)-2\,m=2\,\frac{|q_{B}\,e\,B|}{2\,m}-2\,\frac{4\pi}{2}\,\beta\,|e\,B|^{2}+\order{B^{3}}\,,$ ((48))
[…]
The magnetic polarisability and Landau terms are now isolated as the only remaining field strength dependent terms. This process could be mirrored in a lattice QCD calculation, determining the effective energies in the aligned and anti-aligned case, however this neglects the opportunity to cancel the highly correlated QCD fluctuations contained within the correlation functions of different field strengths and spin alignments.
To optimise this cancellation, we define a spin-field aligned correlator
$G_{\uparrow\uparrow}(B)=G(+s,+B)+G(-s,-B)\,,$ ((49))
where the baryon’s spin is aligned with the magnetic field and a spin-field anti-aligned correlator
$G_{\uparrow\downarrow}(B)=G(+s,-B)+G(-s,+B)\,,$ ((50))
where the spin and field are opposed. To enable the cancellation of QCD fluctuations, we form the correlator ratio
$R(B,t)=\frac{G_{\uparrow\uparrow}(B,t)\,G_{\uparrow\downarrow}(B,t)}{G(0,t)^{2}}\,.$ ((51))
When taking the $\log$ of $R(B,t)$ in determination of the effective energy, the zero-field correlator acts to subtract the mass term, while the numerator product results in the subtraction of the magnetic moment, in an analogous manner to the energy sum of Eq. (48).
As such, we define the magnetic polarisability energy shift $\delta E_{\beta}(B,t)$
$\delta E_{\beta}(B,t)$ $=\frac{1}{2}\frac{1}{\delta t}\lim_{t\rightarrow\infty}\log\Bigl(\frac{R(B,t)}{R(B,t+\delta t)}\Bigr{missing}),$
$=\frac{1}{2}\left[\delta E_{\uparrow\uparrow}(B)+\delta E_{\uparrow\downarrow}(B)\right]-\delta E(0)$
=|$=\frac{\absolutevalue{q_{B}\,e\,B}}{2m}-\frac{4\pi}{2}\beta|B|^{2}+\order*{B^{3}}.$ ) . ((52))
We note that this is the analogue to Eq. (48). The first term, the Landau term, vanishes for neutral baryons allowing for direct access to the magnetic polarisability. For charged baryons, the Landau term must be carefully considered in the fitting process.
Fitting a single-parameter quadratic fit to Eq. IV with the Landau term fully specified as discussed in Sec. VI allows the extraction of the magnetic polarisability.
V Simulation Details
V.1 Gauge ensembles
Table 3: Details of the PACS-CS ensembles used in this work. The lattice spacing of each ensemble is set using the Sommer scale with $r_{0}=0.49\,$fm. In all cases $\kappa_{s}^{\rm sea}=0.13640$ and $\kappa_{s}^{\rm val}=0.13665$ Menadue (2018). $N_{\rm con}$ describes the number of configurations.
$m_{\pi}^{\rm PACS-CS}$(MeV) $\kappa_{u\,d}$ $a\,$(fm) $N_{\rm con}$
701 0.13700 0.1022(15) 399
570 0.13727 0.1009(15) 397
411 0.13754 0.0961(13) 449
296 0.13770 0.0951(13) 399
The four gauge ensembles used in this work are the four heaviest $2+1$-flavour dynamical gauge configurations provided by the PACS-CS collaboration Aoki et al. (2009) through the International Lattice Data Grid (ILDG) Di Renzo (2024). The configurations have a range of degenerate up and down quark masses while the strange quark mass is fixed. The strange quark mass of the ensembles which corresponds to $\kappa_{s}=0.13640$ does not extrapolate to the physical kaon mass Menadue et al. (2012). Use of $\kappa_{s}=0.13665$ for the valence strange quark mass produces the correct value for the kaon mass extrapolated to the physical point Menadue (2018). We note that the mass of the strange quarks in the sea remain at the heavier mass.
Each of the ensembles are a $32^{3}\times 64$ lattice. The gauge action is the Iwasaki gauge action and the clover fermion action with $C_{\rm SW}=1.715$ is the background field corrected clover fermion action Bignell et al. (2019) which is tuned to remove the non-physical magnetic-field induced additive mass renormalisation. Details of the ensembles are summarised in Tab. 3.
As only the valence quarks interact with the background magnetic field the ensembles are electro-quenched. While it is possible to include the background field in the process of generating each gauge field configuration Fiebig et al. (1989), this would require a separate Monte-Carlo simulation at each field strength. Such separate simulations would remove the correlated QCD fluctuations which are otherwise efficiently removed through the ratio in Eq. (51). Without this correlation, a very significant increase in statistics would be required. Instead we preserve the correlations and estimate the small corrections for electro-quenching through chiral effective field theory in the process of extrapolation to the physical point.
V.2 Baryon interpolating fields
The commonly used proton interpolating field in lattice QCD is given by Leinweber et al. (1991)
$\chi_{p}(x)=\epsilon^{abc}\,\left[{u^{a}}^{T}\,C\,\gamma_{5}\,d^{\,b}(x)\right]\,u^{c}(x)\,,$ ((53))
where $C$ is the charge-conjugation matrix. The interpolating fields of the other outer octet baryons may be easily obtained through appropriate substitution of doubly and singly represented quark flavours. This is the form of the interpolating field used throughout this work.
Such interpolating fields, utilised purely with traditional gauge-covariant Gaussian smearing are ineffective at isolating the baryon ground state in a uniform background field Primer et al. (2014); Deshmukh and Tiburzi (2018); Bignell et al. (2020); Bruckmann et al. (2017); Chang et al. (2015). The uniform background field breaks the spatial symmetry and Landau-mode physics presents at both the quark and hadronic levels, and this must be accommodated.
V.3 Quark operators
Asymmetric source and sink operators have been shown to improve the overlap of the lowest energy eigenstates of baryons in a magnetic field Bignell et al. (2020). Following this work we utilise standard Gaussian smearing at the source and a low-lying eigenmode projection at the sink. The low-mode sink projection employs QED+QCD eigenmodes such that the quark propagators are sensitive to the dynamics of nontrivial electric charges in the magnetic field. As the QCD+QED eigenmodes contain the effects of the breaking of spatial symmetry by the uniform field in the $z$-direction, these modes further aid in the isolation of the lowest-energy baryon in the correlation functions. Finally, we include a U(1) projection of the colour-singlet baryon state for charged baryons rather than a traditional Fourier projection. This ensures the isolation of the lowest-Landau-level. Each of these steps is described in detail below.
V.3.1 Link smearing
In constructing the smeared source and sink projections described below, stout link smearing Morningstar and Peardon (2004) is utilised on the spatially-oriented gauge links. 10 smearing sweeps are applied with an isotropic smearing parameter of $\alpha_{stout}=0.1$. These gauge links are used in the process of delta-function source smearing and in the calculation of the sink projection via low-lying eigenmodes of the lattice Laplacian.
V.3.2 Quark propagator source
The quark source is constructed using three-dimensional, gauge-invariant Gaussian smearing Gusken (1990). In the process of smearing, we use stout-links Morningstar and Peardon (2004) as described above.
At all quark masses, $\alpha=0.7$ is used for the Gaussian smearing. The number of gauge-invariant Gaussian smearing sweeps considered is quark mass dependent, with smaller numbers of sweeps associated with heavier quark masses. In tuning the smearing to optimise the onset of early effective-mass plateaus Bignell et al. (2020), 150 through 350 sweeps are utilised.
V.3.3 Boundary conditions
For the calculation of the quark propagators, periodic boundary conditions are used in the spatial dimensions. To avoid signal contamination from the backward propagating states, we use fixed boundary conditions in the temporal direction. The source is then placed at $t=16$, one quarter of the total time-dimension length such that one is always away from the fixed boundary by using the middle part of the lattice time dimension.
V.3.4 Quark propagator sink
The source construction is designed to provide a representation of the QCD interactions with the intent of isolating the QCD ground state. The sink operators are then constructed in such a manner as to encapsulate the quark-level physics of the electromagnetic and QCD interaction. This is done through the eigenmode projection techniques demonstrated in Ref. Bignell et al. (2020), where a comprehensive explanation of the mechanisms may be found.
In particular, the basis of eigenmodes of a fermion operator describing a quark or charged baryon in a constant magnetic field depends only on the lattice Laplacian Bignell et al. (2018). In other words, the Landau modes for a charged Dirac particle in a constant magnetic field $\vec{B}=B\,\hat{z}$ correspond to the eigenmodes of the two-dimensional U(1) gauge-covariant lattice Laplacian. Thus, the fully gauge-covariant sink operator is constructed by first calculating the low-lying eigenmodes of the two-dimensional lattice Laplacian
$\Delta_{\vec{x},\vec{x}^{\prime}}=4\,\delta_{\vec{x},\vec{x}^{\prime}}-\sum_{\mu=1,2}U_{\mu}(\vec{x})\,\delta_{\vec{x}+\hat{\mu},\vec{x}^{\prime}}+U^{\dagger}_{\mu}(\vec{x}-\hat{\mu})\,\delta_{\vec{x}-\hat{\mu},\vec{x}^{\prime}}\,,$ ((54))
where $U_{\mu}(\vec{x})$ are the full SU(3)$\times$×U(1) gauge links described in Sec. III. Again, stout-links are used in constructing the eigenmodes.
Due to the two dimensional nature of the Laplacian, the modes are calculated on each $(z,t)$-slice of the lattice independently. Considering the four-dimensional coordinate space representation of the eigenmode with $\vec{r}=(x,y,z)$
$\expectationvalue{\vec{r},t\,\Big{|}\,\psi_{i,\vec{B}}}=\psi_{i,\vec{B}}(x,y\,|\,z,t)\,,$ ((55))
this may be interpreted as the selection of the two-dimensional eigenmode $\psi_{i,\vec{B}}(x,y)$ in the coordinate space representation on slice $(z,t)$ thus forming the full four-dimensional eigenmode $\psi_{i,\vec{B}}(x,y\,|\,z,t)$.
Through the completeness relation
$1=\sum_{i=1}\,\ket{\psi_{i}}\bra{\psi_{i}}\,,$ ((56))
we form a coordinate-space projection operator
$P_{n}\left(\vec{r},t;\vec{r}^{\prime},t^{\prime}\right)$ $=$ $\sum_{i=1}^{n}\,\expectationvalue{\vec{r},t\,\Big{|}\,\psi_{i,\vec{B}}}\expectationvalue{\psi_{i,\vec{B}}\,\Big{|}\,\vec{r}^{\prime},t^{\prime}}\,$
$=$ $\sum_{i=1}^{n}\,\expectationvalue{x,y\,\Big{|}\,\psi_{i,\vec{B}}}\expectationvalue{\psi_{i,\vec{B}}\,\Big{|}\,x^{\prime},y^{\prime}}\,\delta_{zz^{\prime}}\,\delta_{tt^{\prime}}\,,$
which may then be applied to the quark propagator at the sink. As we are looking for low-energy states, we utilise only the $n$ lowest-lying eigenmodes of the Laplacian. It is shown in Ref. Bignell et al. (2020), that including too few modes results in a noisy hadron correlation function in much the same manner as applying too many sweeps of traditional sink smearing. As such, the number of modes is chosen to be large enough to minimise the noise of the correlation function, but small enough to retain the focus on the aforementioned low-energy physics. The work of Ref. Bignell et al. (2020) found that $n=96$ modes provides balance to these two effects and is what we use here.
V.4 Hadronic projection
The inclusion of the background magnetic field induces a change to the wave function of a charged baryon Roberts et al. (2011). The quark level electromagnetic physics is highlighted by the eigenmode projection at the sink. However, we must also ensure that our operator has the appropriate electromagnetic characteristics on the hadronic level. By projecting final-state charged baryons to the Landau state corresponding to the lowest-lying Landau level, we ensure $n=0$ in Eq. (47).
Due to the colour singlet nature of the baryon, we need only project the eigenmodes of the U(1) Laplacian rather than the full lattice Laplacian used for the sink. Again we follow the formalism of Ref. Bignell et al. (2020).
In the zero-field case, correlators are momentum projected
$G\left(\vec{p},t\right)=\sum_{\vec{x}}\,e^{-i\,\vec{p}\cdot\vec{x}}\,\Braket{\Omega}{T\left\{\chi\left(\vec{x},t\right)\,\bar{\chi}\left(\vec{0},t\right)\right\}}{\Omega},$ ((58))
to $\vec{p}=0$. As neutral baryons do not feel the effects of the background field, these correlation functions are also momentum projected to $\vec{p}=0$.
However, the standard approach of a three-dimensional Fourier projection is not appropriate for a charged baryon when the uniform background magnetic field is present. With a background field present, the baryon’s energy eigenstates are no longer be eigenstates of the $p_{x},p_{y}$ momentum components. Instead, the $x,y$ dependence of the two-point correlator is projected onto the baryon’s lowest Landau level, $\psi_{\vec{B}}\left(x,y\right)$. A Fourier transform of the $z$-coordinate selects a specific value for the $z$-component of momentum
$G\left(p_{z},\vec{B},t\right)=\sum_{x,y,z}\,$ $\psi_{\vec{B}}\left(x,y\right)\,e^{-i\,p_{z}\,z}$
×$\times\Braket{\Omega}{T\left\{\chi\left(\vec{r},t\right)\,\bar{\chi}\left(\vec{0},t\right)\right\}}{\Omega}.$ ((59))
In the infinite-volume continuum limit, the lowest Landau mode has a Gaussian form, $\psi_{\vec{B}}\sim e^{-\absolutevalue{q_{B}e\,B}\,\left(x^{2}+y^{2}\right)/4}$. However, in a finite volume the periodicity of the lattice causes the wave function’s form to be altered Tiburzi and Vayl (2013); Bignell et al. (2020). As such, we instead calculate the lattice Landau eigenmodes using the two-dimensional U(1) lattice Laplacian in an analogous way to Eq. (54) Bignell et al. (2018). Here $U_{\mu}$ contains only the $U(1)$ phases appropriate to the background magnetic field quantised on the lattice.
The correlator projection is then onto the space spanned by the degenerate modes $\psi_{i,\vec{B}}$ associated with the lowest lattice Landau level available to the proton
$G\left(p_{z},\vec{B},t\right)=\sum_{x,y,z}\,\sum_{i=1}^{n}\,$ $\psi_{i,\vec{B}}\left(x,y\right)\,e^{-i\,p_{z}\,z}$
×$\times\Braket{\Omega}{T\left\{\chi\left(\vec{r},t\right)\,\bar{\chi}\left(\vec{0},t\right)\right\}}{\Omega}.$ ((60))
The degeneracy of the lowest-lying Landau mode is governed by the magnetic-field quanta |$\absolutevalue{k_{d}}$ |. Each mode in the finite volume has a degeneracy equal to the magnetic flux quanta, |$\absolutevalue{k_{d}}$ |. Thus, for a baryon with charge $q_{B}$ the degeneracy for the first and second field strengths considered, $k_{d}=\pm 1,\ \pm 2$, are
$n=\left|k_{d}\,\frac{q_{B}}{q_{d}}\right|\,.$ ((61))
For example, the proton’s degeneracy for the first and second field strengths are 3 and 6 respectively.
In evaluating Eq. (60), one can also consider the case of fixing $n=1$, and considering only one of the degenerate eigenmodes through an optimisation procedure. To ensure excellent overlap with the source, one rotates the $U(1)$ eigenmode basis to maximise overlap with the baryon source. Recalling the baryon delta-function source in the $x$-$y$ plane $\rho(x,y)=\delta_{x0}\,\delta_{y0}$, one proceeds to optimise the overlap of the source with the first mode $i=1$ by maximising the value of $\left|\innerproduct{\rho\,}{\,\psi_{i=1,\vec{B}}}\right|^{2}$. An optional phase is then applied such that $\psi_{i=1,\vec{B}}(0,0)$ is purely real at the source point. Illustrations of these eigenmodes are provided in Ref. Bignell et al. (2020).
Ref. Bignell et al. (2020) found that results using the one unrotated degenerate mode where the Landau wave function’s peak is centred on the quark source were almost indistinguishable to those using an optimised linear combination of the degenerate modes. As such, we perform the final-state charged baryon projection using the single degenerate mode peaked at the source. For large field strengths, the probability density of the projection mode has a Gaussian shape. As the field strength decreases the width of the Gaussian increases, making the probability distribution flatter as zero field strength is approached. Landau mode probability densities for charge $|q_{B}|=1$ baryons in fields strengths $|q_{d}|=1$ and $2$ are illustrated in Fig. 2.
Figure 2: Landau mode probability densities for charge $|q_{B}|=1$ baryons at our two lowest field strengths. Recalling $n=|k_{d}\,{q_{B}}/{q_{d}}|$, the $n=3$ (left) and $n=6$ (right) probability densities are illustrated for field strengths $|q_{d}|=1$ and $2$ respectively. The degeneracy of each mode is $n$; the mode best aligned with the fermion source is selected.
This hadronic eigenmode-projected correlator offers superior isolation of the ground state as shown in Ref. Tiburzi and Vayl (2013) and is crucial for the fitting of constant plateaus in the energy shift of Eq. IV herein.
V.5 Statistics
As periodic boundary conditions are used in all four dimensions for the gauge field generation, one can exploit the associated translational invariance of the gauge fields. A quark source can be placed at any position on the lattice and then circularly cycled to the standard source position of $(x,y,z,t)=(1,1,1,16)$. This enables additional sampling of the full gauge field.
Further, the two dimensional nature of the lattice Laplacian operator allows the eigenmodes for the sink projection to be re-used when the gauge field is cycled solely in the time direction. Hence, we increase our statistics on the PACS-CS ensembles by considering four random spatial sources at $t=16$. The gauge field is then circularly cycled in the temporal direction by an eighth of the lattice time extent (eight slices in our case) for each random source. This results in a further increase in statistics by a factor of eight. Together random sources and time-direction cycles increase our statistics by a factor of 32. These multiple samples are binned and averaged as a single configuration estimate in the error analysis.
V.6 Magnetic field
[…]
VI Fitting
As discussed in Sec. IV, we construct the spin-field aligned and anti-aligned correlation functions
$G_{\uparrow\uparrow}(B)$ $=G(+s,+B)+G(-s,-B)\,,$ ((62))
$G_{\uparrow\downarrow}(B)$ $=G(+s,-B)+G(-s,+B)\,,$ ((63))
which are combined in
$R(B,t)=\frac{G_{\uparrow\uparrow}(B,t)\,G_{\uparrow\downarrow}(B,t)}{G(0,t)^{2}}\,,$ ((64))
which aggregates the positive and negative field strengths together to remove the magnetic moment term from the energy expansion of Eq. (47). Due to this aggregation, any reference to the magnetic field strength from this point refers to the aggregated positive field strength.
Recalling Eq. IV, we then construct the magnetic polarisability energy shift
$\delta E_{\beta}(B,t)$ $=\frac{1}{2}\frac{1}{\delta t}\lim_{t\rightarrow\infty}\log\Bigl(\frac{R(B,t)}{R(B,t+\delta t)}\Bigr{missing}),$
=|$=\frac{\absolutevalue{q_{B}\,e\,B}}{2m}-\frac{4\pi}{2}\beta|B|^{2}+\order*{B^{3}}.$ ) . ((65))
Fitting to the above ratio in the long time limit and repeating this as a function of our two positive field strengths provides access to the magnetic polarisability.
In practice it is simple to fit in terms of the field strength quanta $k_{d}$. Using the quantisation condition of Eq. (46)
$e\,B=\frac{2\,\pi}{N_{x}\,N_{y}\,a^{2}}\frac{1}{q_{d}}\,k_{d},$ ((66))
we substitute for $e\,B$ in Eq. VI to re-write $\delta E_{\beta}$ in terms of the field strength quanta
$\delta E_{\beta}(k_{d})$ $=\frac{q_{B}}{2m}\absolutevalue{\frac{2\,\pi}{N_{x}\,N_{y}\,a^{2}}\frac{1}{q_{d}}\,k_{d}}-\frac{4\pi}{2}\beta\left(\frac{2\,\pi}{e\,N_{x}\,N_{y}\,a^{2}}\frac{1}{q_{d}}\,k_{d}\right)^{2},$
$=\frac{2\pi}{N_{x}\,N_{y}\,a^{2}}\frac{1}{2\,m}\absolutevalue{\frac{q_{B}}{q_{d}}\,k_{d}}-\frac{1}{2\,\alpha}\left[\frac{2\pi}{N_{x}\,N_{y}\,a^{2}}\right]^{2}\frac{1}{q_{d}^{2}}\,\beta\,k_{d}^{2},$ ((67))
$\equiv L(k_{d},m)+C\,\beta\,k_{d}^{2}\,.$ ((68))
It is important to note the distinction between $q_{B}$ which is the charge of the baryon in question and $q_{d}$, the down quark charge introduced in the magnetic field quantisation condition. Here $\alpha=\frac{e^{2}}{4\pi}$ is the fine structure constant. For convenient future reference, we have defined $L(k_{d},m)$ the Landau term and $C$ the remaining coefficient to the magnetic polarisability term.
Previous fitting procedures Bignell et al. (2018, 2020) for the magnetic polarisability in the background field method followed the following steps
  • Consider all field strengths under investigation and select a common fit window for the magnetic polarisability energy shift. Fit the energy shift to determine $\delta E_{\beta}(k_{d})$.
  • Construct $\delta E_{\beta}(k_{d})-L(k_{d},m)$ by subtracting the known Landau term at each field strength using the best estimate for the baryon mass on that ensemble.
  • Fit $\left[\delta E_{\beta}(k_{d})-L(k_{d},m)\right]/C$ as a function of the integer $k_{d}^{2}$ to obtain the magnetic polarisability.
Specifically, fits for $\delta E_{\beta}(k_{d})$ were judged by the full covariance matrix $\chi^{2}_{\rm dof}$ coefficient, estimated by the jackknife method Efron (1979). We required $\chi^{2}_{\rm dof}<1.2$ at both field strengths with the start and end points of the fit window being common for each field strength. These criteria are designed to maximise the cancellation of correlated QCD fluctuations.
Figure 3: A plateau fit for a proton at $m_{\pi}=296\,$MeV. This effective energy plateau fit to $\delta E_{\beta}(B,t)$ is an example of a fit that looks acceptable, but is in fact fit too early due to underlying excited state contamination.
An example of such a fit is shown in Fig. 3.
VI.1 Fitting issues
This method presents a couple of issues. It makes the assumption that once the $\delta E_{\beta}(k_{d},t)$ energy shift plateaus, the underlying correlation functions used to construct the ratio are also independently exhibiting plateau-like behaviour. That is an assumption that a plateau in $\delta E_{\beta}(k_{d},t)$ implies single state isolation of all underlying correlation functions. We have seen that this is not necessarily the case.
Consider the plateau plot for the proton in Fig. 3 at $\kappa=0.13770$, $m_{\pi}=296\,$MeV. The selected fit window [21, 27] appears very reasonable. The $\chi^{2}_{\rm dof}$ values are both acceptable, the energy shift appears plateau-like at each field strength and we don’t appear to be fitting significant noise.
However, we need to check that the underlying correlators that form the ratio $R(k_{d},t)$ of Eq. (51) have all reached single state isolation.
[…]
The general form of these correlators is
$G(k_{d},t)=\sum_{\alpha}e^{-E_{\alpha}t}\lambda^{\alpha}\bar{\lambda}^{\alpha}\,,$ ((69))
where $E_{\alpha}$ are the energies of this constructed state and $\lambda^{\alpha},\bar{\lambda}^{\alpha}$ are the couplings of the baryon interpolating fields to the baryon states at the sink and source respectively. Due to the asymmetric source and sink operators used here, the couplings are not adjoint. Taking the log
$\log(G_{A}(B,t))\stackrel{{\scriptstyle t\rightarrow\infty}}{{=}}\log(\lambda\bar{\lambda^{\prime}})-E_{0}\,t\,,$ ((70))
which reaches a linear form in the long time limit once single-state isolation has been reached. As such, the $\chi^{2}_{\rm dof}$ of a linear fit to $\log G$ is an excellent metric for single state isolation of the underlying correlator.
To provide an easy interface to the vast number of $\chi^{2}_{\rm dof}$ values which need to be checked, we produce a series of heat maps. The heat maps corresponding to the underlying correlators of the proton magnetic polarisability are shown in Fig. 4.
Each plot corresponds to a different underlying correlator. We must consider the zero-field correlator in addition to the aligned $G_{\uparrow\uparrow}$ and anti-aligned $G_{\uparrow\downarrow}$ correlators at each field strength. In each case we consider all windows which may be of interest. The heat maps clearly show that $t_{s}=21$, the starting time slice chosen in Fig. 3 is too early.
Noting that the $\chi^{2}_{\rm dof}$ has a distribution, our absolute criteria for the choice of a starting time slice is the first three fit windows must satisfy the criteria $\chi^{2}_{\rm dof}<2.5$. We regard this criteria as a cut to ensure the fit windows considered in the averaging procedure described below are relevant and do not adversely affect the final result. At the same time, it is desirable to consider several meritorious fit windows such that an average weighted by the $\chi^{2}_{\rm dof}$ may be considered. The details of the weighted averaging are provided in subsection VI.3.
The criteria $\chi^{2}_{\rm dof}<2.5$ corresponds to the three leftmost boxes in a row being blue, green, or yellow and this corresponds to plateau-like behaviour across 5 time slices. Earliest possible starting time slices in each case according to the criteria have been highlighted with green text in the column denoting $t_{s}$.
Based on this, we see that we cannot begin fitting until time slice $t=24$, much later than $t=21$ which we selected based only on the plateau behaviour of the polarisability energy shift. This highlights the importance of checking the single state isolation of the underlying correlators. Pushing the fit later here changes the result significantly. We do note that in many cases, the polarisability energy shift does not plateau until the underlying correlators have reached single-state isolation, however it is important to check due to cases such as the proton presented here.
Finally, we do not require the fit for $\log G$ to maintain a good quality of fit indefinitely as the correlation function can suffer a loss of signal. Instead, we use the measure as a metric for excited state contamination. We need only determine when $\log G$ becomes linear. As an example, consider the spin-field aligned, $k_{d}=2$, $\chi^{2}_{\rm dof}$ heat map for the neutron at $\kappa=0.13700$, $m_{\pi}=701\,$MeV shown in Fig. 5. Here we see two clear regions. The criteria is satisfied for fit windows $[t_{s},t_{f}]=[22,28],\ [23,28]$ and 24,28]24,28]24 , 28 ]. Then again at $t_{s}=28$ onwards. In this case, we take the earliest start point to be $t_{s}=22$, as the $\chi^{2}_{\rm dof}$ suggests clear single state isolation over 7 time slices. This scale is long enough to eliminate the possibility of a short-lived false plateau associated with an excited state. Further inspection reveals the source of the red and black blocks in the heat map commencing at $t_{f}=29$ is noise due to a loss of signal.
Figure 5: $\chi^{2}_{\rm dof}$ heat map for fits to $\log G$ for all fit windows $t_{s}\to t_{f}$ of the spin-field aligned correlator at the second field strength. Here we see the behaviour exhibited in some cases in which the quality of $\log G$ fit is good in two places and unsatisfactory in between. This example is a neutron at $\kappa=0.13700$, $m_{\pi}=701$ MeV. Further inspection reveals the red quadrant is associated with a loss of signal in the correlator.
VI.2 Fit window dependence
With confidence in hand that we are choosing fit windows with reasonable suppression of excited state contamination, we still need to select a fit window. This time we consider the proton at $\kappa=0.13727$, $m_{\pi}=570\,$MeV.
The start point criteria derived from the heat maps as described in the previous section requires $t_{s}=29$ in this case. Based purely on the $\chi^{2}_{\rm dof}$ values for the two field strengths, any fit window between $t=29$ and $t=35$ is reasonable. In many cases, such a flexibility in the choice of window produces minimal variation in the resulting value. However, this is not the case here. A summary of the polarisability obtained from various acceptable fit windows are summarised in Tab. 4. Here we see some variation in the resulting magnetic polarisability. This presents a problem as we desire a systematic and consistent scheme through which we can extract the magnetic polarisability. To resolve this issue, we employ a weighted averaging method Beane et al. (2021) to systematically extract a consistent result from a given set of fitting regions.
Table 4: Dependence of the magnetic polarisability of the proton, $\beta_{p}$, on the choice of fit window at $m_{\pi}=570\,$MeV. Magnetic polarisability values in are units of ×$\times 10^{-4}\,$fm3.
$t_{s}$ $t_{f}$ $\beta_{p}$ $\chi^{2}_{{\rm dof},k_{d}=1}$ $\chi^{2}_{{\rm dof},k_{d}=2}$
29 32 2.41(19) 0.87 0.42
29 34 2.35(19) 0.58 0.56
30 34 2.22(25) 0.68 0.70
31 34 2.01(33) 0.41 0.46
32 34 1.78(46) 0.21 0.47
VI.3 Weighted Averaging
Our weighted averaging method is based on that described in the Appendix of Ref. Beane et al. (2021). Within a region $t\in[t_{\rm min},t_{\rm max}]$, all eligible fits are weighted based on their uncertainty, $\chi^{2}$, and number of degrees of freedom. We choose $t_{\rm min}$ based on the heat maps as discussed above. $t_{\rm max}$ is chosen simply as the final time slice before signal is clearly lost to noise. This corresponds to $t_{\rm max}=27$ in Fig. 3.
We consider a specific fit window $[t_{s},t_{f}]$ to be eligible if the window has minimum length three ($t_{f}\geq t_{s}+2$) and $t_{f}=t_{\rm max}$. Fixing $t_{f}=t_{\rm max}$ ensures the collection of possible fit windows is sampled equitably. Consideration of every possible end point for a fit starting at a given time slice would favour fit windows commencing at early time slices; they would contribute on more occasions to the weighted average than a fit that starts later. Thus the criteria $t_{f}=t_{\rm max}$ ensures each starting time slice is considered once.
For example in a fitting range $t\in[20,24]$ where all starting points are equally favourable, the possible windows are $[20,22],[20,23],[20,24],[21,23],[21,24],[22,24]$. We clearly see that windows with $t_{s}=20$ would be favoured in the calculation due simply to occurring earlier. Fixing $t_{f}=t_{\rm max}$ provides $[20,24],[21,24]$ and $[22,24]$ which equitably samples possible $t_{s}$ values. This is important as the value of a fit is mostly determined by the first few data points contained in a fit where the uncertainties are small. Over counting early fits is likely to lead to systematic errors.
With the candidate fit windows determined, the $i$th window of $N$ candidates is assigned a weight according to
$w_{i}=\frac{1}{\mathcal{Z}}\frac{p_{i}}{(\delta E_{i})^{2}}\,,$ ((71))
where
$\mathcal{Z}=\sum_{i=1}^{N}\frac{p_{i}}{(\delta E_{i})^{2}}\,.$ ((72))
Here, $\delta E_{i}$ is the uncertainty of the $i$th fit, and $p_{i}$ is the $p$-value of the fit. The $p$-value is the probability of a given fit occurring in a $\chi^{2}$ distribution. It is most easily calculated using the gamma distribution and gamma function in the following manner
$p_{i}=\frac{\Gamma(N_{\rm dof}/2,\chi^{2}/2)}{\Gamma(N_{\rm dof}/2)}\,.$ ((73))
Dividing the degrees of freedom and $\chi^{2}$ of the fit by two and normalising the gamma distribution in the numerator with the gamma function in the denominator causes the gamma distribution to rescale precisely to the $\chi^{2}$ distribution, hence producing the probability of any given fit occurring based on its $\chi^{2}$ and number of degrees of freedom.
With these weights, the average effective energy is
$E=\sum_{i}^{n}w_{i}\,E_{i}\,,$ ((74))
and the statistical error
$(\delta E)^{2}=\sum_{i}^{n}w_{i}(\delta E_{i})^{2}\,.$ ((75))
We now apply this method to the proton correlators considered in Tab. 4. We assign a start point $t_{\rm min}=29$ based on the heat map of the underlying correlators and choose $t_{\rm max}=34$. Beyond that, all signal has been lost to noise.
The weights for each window are shown in Tab. 5. We can see that the later windows, which are more susceptible to noise due to their larger uncertainties are suppressed.
Table 5: Fit weights of Eq. (71) for the proton at $m_{\pi}=570\,$MeV. Strength is distributed among the candidate fit windows with later fit windows suppressed. Magnetic polarisability values are in units of ×$\times 10^{-4}\,$fm3.
$t_{s}$ $t_{f}$ $\beta_{p}$ $\chi^{2}_{{\rm dof},k_{d}=1}$ $\chi^{2}_{{\rm dof},k_{d}=2}$ $w_{k_{d}=1}$ $w_{k_{d}=2}$
28 34 2.35(15) 0.51 0.77 0.23 0.33
29 34 2.35(19) 0.58 0.56 0.28 0.23
30 34 2.22(25) 0.68 0.70 0.31 0.27
31 34 2.01(33) 0.41 0.46 0.13 0.11
32 34 1.78(46) 0.21 0.47 0.04 0.06
While our procedure assigns the same $t_{\rm min}$ and $t_{\rm max}$ to both $k_{d}$ values governing the field strength, we fit and weight the two field strengths independently. It is not uncommon for their correlators to have significantly different $\chi^{2}_{\rm dof}$ values in spite of their correlated nature. We also explored the possibility of forcing common weights for the field strengths to maximise the opportunity for cancellation of correlated fluctuations. However, we observed this approach to have minimal effect on the final uncertainty. As a result, we fit and weight the two field strengths independently.
Having defined values for $\delta E_{\beta}(k_{d})$ at $k_{d}=1,2$, we turn our attention to the Landau term $L(k_{d},m)$ of Eq. (68), which depends on the baryon mass. This mass is also determined through the weighted average approach.
The magnetic polarisability is determined in a single parameter fit of
$\frac{\delta E_{\beta}(k_{d})-L(k_{d},m)}{C}=\beta\,k_{d}^{2}\,.$ ((76))
Examples of such fits are shown in Fig. 6.
Figure 6: Fits to Eq. (76) as a function of $k_{d}$. These representative examples are at $\kappa=13754$, $m_{\pi}=411\,$MeV.
To determine the uncertainty in the polarisability, a jackknife error estimate is performed. A second-order jackknife is used to obtain uncertainties on the correlation functions and correlation function ratios. Uncertainties for fit values such as the magnetic polarisability energy shift for each field strength are obtained from individual first-order jackknife sub-ensembles. Finally, the fit of the energy shifts as a function of field strength is repeated on each jackknife sub-ensemble and the error on the ensemble average calculated as the jackknife error.
In this particular case, we obtain $\beta_{p}=2.25(21)\times 10^{-4}\,$fm3 for the proton at $\kappa=0.13727$, $m_{\pi}=570\,$MeV. This process is repeated for every baryon at every quark mass.
VII Lattice Results
Magnetic polarisability values are presented in Tab. 6. and illustrated in Fig. 7 in the context of the constituent quark model of Sec. II. One observes a discrepancy in magnitude between the quark model predictions and the lattice results and we will address this below.
However, a few observations of the lattice results are worthy of note. The uncertainties in the lattice results are sufficiently small to reveal interesting structure in the octet baryon polarisabilities. The $\Xi^{0}$ stands out in magnitude in accord with quark model expectations. This is followed by a cluster of $n$, $p$ and $\Sigma^{+}$ magnetic polarisabilities of similar magnitude, again in accord with the quark model. Finally, the negatively charged $\Xi^{-}$ and $\Sigma^{-}$ baryons have very small polarisabilities as the quark model predicts.
Table 6: Magnetic polarisability results for the outer octet baryons on the PACS-CS ensembles. Polarisability values are in units of ×$\times 10^{-4}\,$fm3, and pion masses are in the Sommer scheme.
$\kappa$ $m_{\pi}/$MeV $p$ $n$ $\Sigma^{+}$ $\Xi^{0}$ $\Sigma^{-}$ $\Xi^{-}$
0.13700 623 2.32(14) $2.390(98)$ 2.37(13) $2.50(10)$ $-0.06(16)$ $-0.03(12)$
0.13727 515 2.25(21) $2.14(13)$ 2.29(18) $2.364(89)$ $-0.130(73)$ $-0.12(11)$
0.13754 390 1.64(12) $1.66(12)$ 1.85(12) $2.36(17)$ $-0.05(16)$ $-0.01(14)$
0.13770 280 1.41(25) $1.40(22)$ 1.57(12) $2.81(27)$ $0.23(14)$ $-0.03(10)$
Figure 7: Uncorrected (a) and corrected (b) quark model predictions for the magnetic polarisability on the PACS-CS ensembles are displayed as interpolated curves. The lattice data from Tab. 6 are plotted as points on both figures. The legend is common to both plots and ordered to match the vertical ordering at the first dashed line. Dashed lines represent the pion masses of the PACS-CS ensembles used in this work.
VII.1 Corrected Quark Model
To address the magnitude difference between our lattice QCD results and the quark model, we consider a simple correction to the quark model. Recalling Eq. II.1
[…]
$\equiv\beta_{1}-\beta_{2}\,,$ ((77))
there are two contributions to the magnetic polarisability. To retain the negative contribution of $\beta_{2}$, we include two positive fit parameters $a_{1},\,a_{2}$
$\beta=a_{1}\,\beta_{1}-a_{2}\,\beta_{2}\,,$ ((78))
and correct the model by fitting these two parameters to describe the magnetic polarisabilities of all six octet baryons at each of the four quark masses considered in lattice QCD. A simple least squares minimisation produces the fit parameters
$a_{1}=0.417,\quad a_{2}=0.540\,.$ ((79))
We see an approximate reduction in both terms of a half. This corrected model is illustrated in Fig. 7 beside the original uncorrected model. We see that the quark model now broadly agrees in a qualitative manner with the lattice values and provides a detailed understanding of the physics underpinning the magnetic polarisabilities of octet baryons. Given the simplicity of the constituent quark model, the quality of, and insight provided by its predictions are significant.
We see excellent agreement for the $\Xi^{0}\,$. We also see the clear difference in polarisability of the $\Xi^{-}\,$and $\Sigma^{-}\,$as predicted by the quark model, though we lack sufficient precision to make a statement about the sign of those magnetic polarisabilities.
The ordering of baryons in the lattice data is very interesting. Correlated differences between the proton, neutron and $\Sigma^{+}\,$cannot order them uniquely at one sigma. The ordering changes for different quark masses.
We also highlight the quark mass dependence of the proton, neutron and $\Sigma^{+}\,$in the lattice results. We see a clear trend downwards as we approach the physical point. However, this trend is not captured well by the quark model.
We now proceed to examine the implications of the chiral physics which dominate the interactions near the physical point and connect these lattice results with experiment.
VIII Chiral extrapolation
VIII.1 Formalism
Figure 8: Two-photon interactions relevant to the baryon magnetic polarisability. In coupling the photons to the intermediate meson, these diagrams generate the most important leading-nonanalytic behaviour in the quark mass dependence of the magnetic polarisability of a baryon. Intermediate baryon states include those degenerate with the initial baryon (left) and those where the intermediate state has a mass different from the initial baryon (right).
To complete our analysis, we now connect our lattice QCD results to the physical world via chiral extrapolation. In doing so, we draw on chiral effective field theory ($\chi$EFT) to inform corrections for systematic uncertainties associated with the finite-volume of the lattice and electro-quenching of the quark sea. This analysis follows the formalism established in Refs. Hall et al. (2014) and Bignell et al. (2020), and extends the formalism to address all six of the outer octet baryons considered herein.
As our lattice QCD results can be described well by a fit linear in $m_{\pi}^{2}$, we consider the following chiral expansion for the magnetic polarisability
$\beta^{B}(m_{\pi}^{2})=a_{0}^{\Lambda}$ $+a_{2}^{\Lambda}\,m_{\pi}^{2}$
$+\sum_{M}\,\beta^{MB}(m_{\pi}^{2},\Lambda)$
$+\sum_{M,B^{\prime}}\,\beta^{MB^{\prime}}(m_{\pi}^{2},\Lambda)+\order{m_{\pi}^{3}},$ ((80))
where $a_{0}^{\Lambda}$, and $a_{2}^{\Lambda}$ are residual series coefficients which we constrain using volume-corrected lattice QCD results. $\Lambda$ is the renormalisation scale and $\beta^{MB}$ and $\beta^{MB^{\prime}}$ are the leading-order loop contributions to the magnetic polarisability from the diagrams of Fig. 8. The scale-dependent residual-series coefficients are combined with the analytic scale-dependent terms of the loop contributions to recover the renormalised series expansion Young et al. (2003); Hall et al. (2010, 2012, 2013).
The loop contributions have integral forms in the heavy-baryon approximation given by
$\beta^{MB}(m_{\pi}^{2},\Lambda)=\frac{e^{2}}{4\pi}\frac{1}{288\,\pi^{3}\,f^{2}_{\pi}}\,\chi_{MB}\int d^{3}k\,\frac{\vec{k}^{2}\,u^{2}(k,\Lambda)}{\omega^{6}_{\vec{k}}}\,,$ ((81))
and for $\Delta M=m_{B^{\prime}}-m_{B}$
$\beta^{MB^{\prime}}(m_{\pi}^{2},\Lambda)=\frac{e^{2}}{4\pi}\frac{1}{288\,\pi^{3}\,f^{2}_{\pi}}\,\chi_{MB^{\prime}}\int d^{3}k\,{\vec{k}^{2}\,u^{2}(k,\Lambda)}\,\frac{\omega^{2}_{\vec{k}}\,\Delta M\left(3\,\omega_{\vec{k}}+\Delta M\right)+\vec{k}^{2}\left(8\,\omega^{2}_{\vec{k}}+9\,\omega_{\vec{k}}\,\Delta M+3\,(\Delta M)^{2}\right)}{8\,\omega^{5}_{\vec{k}}\left(\omega_{\vec{k}}+\Delta M\right)}\,,$ ((82))
where $\omega_{\vec{k}}=\sqrt{\vec{k}^{2}+m_{M}^{2}}$ is the energy carried by the meson $M$ with three-momentum $\vec{k}$, $f_{\pi}=92.4\,$MeV is the pion decay constant and
$u(k,\Lambda)=\frac{1}{(1+\vec{k}^{2}/\Lambda^{2})^{2}},$ ((83))
is a dipole regulator which ensures only soft momenta flow through the effective-field theory degrees of freedom. Eq. (81) generates the leading nonanalytic contribution to the magnetic polarisability proportional to $1/m_{\pi}$. It is through these known contributions that we can correct for the electro-quenched nature of the calculation and estimate finite-volume corrections to the lattice results.
For each of the baryons in this work, $p$, $n$, $\Sigma^{\pm}$, $\Xi^{0,-}$, we consider transitions to all possible octet and decuplet baryons. These transitions involve the mesons $\pi^{\pm,0}$, $K^{\pm,0}$, $\overline{K}^{0}$, $\eta$, and $\eta^{\prime}$. We note that all transitions featuring a neutral meson in the intermediate state vanish in the leading order terms of full QCD, but not in the electro-quenched theory where sea-quark charges are effectively zero.
To correct for the electro-quenching present in the lattice results we must first determine the contributions $\chi_{MB}$ and $\chi_{MB^{\prime}}$ associated with $\beta^{MB^{\prime}}$ and $\beta^{MB^{\prime}}$ respectively. These contributions are derived from the interactions in Fig. 8 with reference to the quark flow diagrams of Fig. 9. There are two cases to examine; one where the quark flow is fully connected and one with a disconnected sea-quark-loop contribution.
Figure 9: The fully connected quark-flow diagram (a) features only valence quarks. The quark-flow disconnected diagram (b) contains a sea-quark-loop contribution to a baryon transition to an intermediate meson-baryon state. In this case, we have a proton, $p$, transitioning to an intermediate $\pi^{+}n$ intermediate state and it is straight forward to extend these quark flow diagrams to consider any of the outer octet baryons with doubly and singly represented quark flavours.
In each case we must attach two photons to the meson that is formed in the interaction. In the connected case, all quarks must be connected to valence quarks, however in the disconnected case we have the option of connecting one or two photons to a disconnected sea quark. As such, we have three types of interactions. We call these valence-valence, valence-sea and sea-sea. In equations, these are abbreviated to their initials.
In the lattice results, only the valence-valence contributions are present. The sea quarks are electrically neutral and diagrams where a photon couples to a sea quark vanish. As such, some contributions to the magnetic polarisability are not present and some contributions which should cancel out in summing the contributions of connected and disconnected quark flows do not do so.
As the chiral coefficients are analytically known, we can proceed by fitting the lattice QCD results in the electro-quenched effective field theory and then explicitly add the missing disconnected sea-quark contributions. Each contribution contributes proportionally with the charge of the interacting quarks, so in the example $p\rightarrow n\,\pi^{+}$ shown in Fig. 9 (right) we have
$\chi_{v-v}$ $\propto q_{u}^{2},$ ((84))
$\chi_{v-s}$ $\propto 2q_{u}\,q_{\bar{d}},$ ((85))
$\chi_{s-s}$ $\propto q_{\bar{d}}^{2},$ ((86))
where the factor of two comes from the two possible orderings of the photon attachment.
The coupling strengths may be obtained in a partially-quenched chiral perturbation theory scheme Leinweber (2004) where the flavour of the disconnected sea quark is relabelled to the missing SU(3) quark flavour. In this way the quark disconnected contribution can be isolated in terms of known meson-baryon dressing coefficients. For the case under consideration here, that is the strange quark. In this case the intermediate state would become $\Lambda/\Sigma^{0}$ + $K^{+}$, hence the sea-sea contribution becomes
$\chi_{s-s}$ $\propto q_{\bar{d}}^{2}\,(\chi_{K^{+}\Sigma^{0}}^{2}+\chi_{K^{+}\Lambda}^{2}),$
$=q_{\bar{d}}^{2}\,\bigl{(}(D-F)^{2}+\frac{1}{3}(D+3F)^{2}\bigr{)}\,,$ ((87))
where $F$ and $D$ are the standard axial coupling constants. The valence-sea contributions are obtained equivalently by simply adjusting the charge coefficients and counting the two possible orderings for the photon couplings
$\chi_{v-s}$ $\propto 2\,q_{u}\,q_{\bar{d}}\,(\chi_{K^{+}\Sigma^{0}}^{2}+\chi_{K^{+}\Lambda}^{2}),$
$=2\,q_{u}\,q_{\bar{d}}\,\bigl{(}(D-F)^{2}+\frac{1}{3}(D+3F)^{2}\bigr{)}\,.$ ((88))
With the disconnected quark-flow contributions determined, the valence-valence contribution is obtained by subtracting the valence-sea and sea-sea contributions from the full QCD contribution. In some cases the full QCD contribution may be zero leading to equal and opposite contributions from the connected and disconnected quark flow contributions. However, in the case under consideration here, the $p\to n\pi^{+}$ transition makes a contribution proportional to $\chi_{\pi^{+}n}^{2}$ such that
$\chi_{v-v}$ $\propto q_{\pi^{+}}^{2}\,\chi_{\pi^{+}n}^{2}-(2\,q_{u}\,q_{\bar{d}}+q_{\bar{d}}^{2})\,(\chi_{K^{+}\Sigma^{0}}^{2}+\chi_{K^{+}\Lambda}^{2}),$
$=(2D+F)^{2}$
$\quad-(2\,q_{u}\,q_{\bar{d}}+q_{\bar{d}}^{2})\,\bigl{(}(D-F)^{2}+\frac{1}{3}(D+3F)^{2}\bigr{)}\,.$ ((89))
This process is repeated for all possible transitions for each of the outer octet baryons. All relevant chiral coefficients and associated contributions are summarised in Appendix A.
We note that we have not included transitions to an intermediate $\bar{s}s$ pseudoscalar meson. While it does not contribute to full QCD processes, its consideration can in principle lead to a contribution in the process of separating valence and sea contributions. However, the mass of the $\bar{s}s$ pseudoscalar meson is large at approximately $\sqrt{2m_{K}^{2}-m_{\pi}^{2}}\simeq 685\,$MeV. In the finite-range regularisation used here, the contributions from such large-mass mesons are naturally suppressed, such that the partially-quenched correction is negligible.
It’s interesting to note that transitions involving $\pi^{0}$, $\eta$ and $\eta^{\prime}$ intermediate mesons do not contribute to this analysis. In full QCD these mesons are electrically neutral and do not generate a contribution to the leading loop integrals. In partially quenched QCD, their composition of $\bar{u}u$, $\bar{d}d$, and $\bar{s}s$ matched flavours means they are not relevant to the disconnected flow calculations which by definition include a third quark flavour that differs from the quark flavours present in the outer octet baryons under consideration. As such, subtleties associated with SU(3) flavour symmetry breaking in the $\pi^{0}$, $\eta$ and $\eta^{\prime}$ masses do not enter Leinweber and Thomas (2024).
VIII.2 Implementation
In the process of addressing finite-volume corrections and electro quenching, we proceed as follows. We first leverage the leading loop integrals to inform the size of finite volume effects. The finite volume of the lattice requires infinite-volume integrals over momenta such as those in Eq. (81) and Eq. (82) to be calculated as sums. As such, the finite volume effects associated with the effective field theory are given by the difference between the continuous infinite-volume integrals and the discrete sums. Hence, the finite-volume corrected (FVC) valence-valence lattice polarisability is
$\beta^{\rm FVC}_{v-v}(m_{\pi}^{2})=$ $\beta^{\rm lattice}_{v-v}(m_{\pi}^{2})$
$+$ $\sum_{MB}\,\Big{(}\beta^{MB}_{\rm integ}(m_{\pi}^{2},\Lambda^{FV})-\beta^{MB}_{\rm sum}(m_{\pi}^{2},\Lambda^{FV})\Big{)}$
$+$ $\sum_{MB^{\prime}}\Big{(}\beta^{MB^{\prime}}_{\rm integ}(m_{\pi}^{2},\Lambda^{FV})-\beta^{MB^{\prime}}_{\rm sum}(m_{\pi}^{2},\Lambda^{FV})\Big{)}.$ ((90))
where the idea is to subtract the finite-volume sum (sum) contained within the lattice results and replace it with the infinite-volume integral (integ). Here, the regulator parameter $\Lambda^{FV}$ is selected to be different from that of the main expression as one needs to avoid a collision between infrared and ultraviolet effects Hall et al. (2014). We consider, $\Lambda^{FV}=2.0\,$GeV as in Ref. Hall et al. (2014).
The residual series coefficients $a_{0}(\Lambda)$,$a_{2}(\Lambda)$ are now obtained by fitting to the finite volume corrected, valence-valence lattice results according to
$\beta^{\rm FVC}_{v-v}(m_{\pi}^{2})=a_{0}^{\Lambda}$ $+a_{2}^{\Lambda}\,m_{\pi}^{2}$
$+\sum_{M}\,\beta^{MB}_{v-v}(m_{\pi}^{2},\Lambda)$
$+\sum_{M,B^{\prime}}\,\beta^{MB^{\prime}}_{v-v}(m_{\pi}^{2},\Lambda)+\order{m_{\pi}^{3}},$ ((91))
where the subscripts ${v-v}$ indicate the use of the valence-valence chiral coefficients in place of the full QCD coefficients $\chi_{MB}$ and $\chi_{MB^{\prime}}$ in Eqs. Eq. (81) and Eq. (82).
This time we take the phenomenologically motivated value from the induced pseudoscalar form factor of the nucleon with $\Lambda=0.8\,$GeV for the regulator parameter Wang et al. (2009a); Young et al. (2002); Leinweber et al. (2005, 2006); Wang et al. (2009b). Such a value defines a pion cloud contribution to masses Young et al. (2002), magnetic moments Leinweber et al. (2005), and charge radii Wang et al. (2009a) allowing for the correction of the sea-quark-loop contributions to the pion cloud which play a significant role, especially at small pion masses. At this regulator mass, the nucleon core contribution governed by the residual series coefficients is insensitive to sea-quark-loop contributions.
With the values for $a_{0}^{\Lambda}$ and $a_{2}^{\Lambda}$ determined in the fit of Eq. (91), the electro-quenching correction is incorporated by simply replacing the valence-valence coefficients by the full-QCD coefficients, i.e. by simply evaluating the full expression for the magnetic polarisability in Eq. (80). The resulting values for the octet-baryon magnetic polarisabilities at the physical pion mass $m^{\rm phys}_{\pi}=140\,$MeV are given in Tab. 7.
Systematic uncertainties associated with the chiral extrapolation are estimated through variation of the regulator parameter. The parameter is varied over the broad range $0.6\,\text{GeV}\leq\Lambda\leq 1.0\,$GeV to allow the estimation of the uncertainty associated with the higher terms of the chiral expansion. This systematic uncertainty is also included in Tab. 7.
Table 7: Magnetic polarisability values for the outer octet baryons in full QCD at the physical pion mass, $m^{\rm phys}_{\pi}=0.140\,$GeV, in infinite volume. The values include both finite-volume and electro-quenching corrections as described in the text. All values are in units of ×$\times 10^{-4}\,$fm3. Uncertainties include statistical simulation uncertainties and systematic uncertainties associated with the chiral extrapolation. These are combined in quadrature in the final column.
Uncertainties
Baryon $\beta$ Statistical Systematic Combined
$p$ 2.10 0.17 0.16 0.23
$n$ 2.11 0.15 0.16 0.22
$\Sigma^{+}$ 1.83 0.12 0.06 0.13
$\Sigma^{-}$ 0.51 0.15 0.09 0.17
$\Xi^{0}$ 2.58 0.18 0.04 0.19
$\Xi^{-}$ 0.11 0.11 0.03 0.11
(a) Proton extrapolation.
(b) Neutron extrapolation.
(c) $\Sigma^{+}\,$extrapolation.
(d) $\Sigma^{-}\,$extrapolation.
(e) $\Xi^{0}\,$extrapolation.
(f) $\Xi^{-}\,$extrapolation.
Figure 10: Chiral extrapolation of the finite-volume- and electro-quenched-corrected (FV & EQ Corr.) octet-baryon magnetic polarisabilities to the physical pion mass indicated by the vertical dashed line. Uncorrected lattice points (Lattice Points) are horizontally offset from the corrected points. Extrapolation curves for future larger finite volumes (FV: $\cdots$ fm) are shown to illustrate the requirements for observing chiral curvature in full QCD. The curves are shown for $m_{\pi}\,L>$ 3, where $L$ is the spatial lattice length. Finally, the infinite-volume extrapolation (Inf. Vol.) relevant to nature is illustrated.
It’s interesting to note that one can also use Eq. VIII.2 to explore the chiral curvature to be seen in future lattice QCD calculations. This time, one subtracts the infinite-volume integral, and adds a finite volume sum appropriate to the future lattice simulation volume. This idea is explored as we present the chiral extrapolations.
Extrapolations of the octet-baryon magnetic polarisabilities to the physical point are presented in Fig. 10. The $p$, $n$, and $\Sigma^{+}$ baryons have very similar profiles as predicted by the simple quark model of Sec. II. Recall the $\Xi^{0}$ baryon is unique due to its small octet-decuplet mass splitting and containing only one $u$ quark which contributes well to the transition term of the model while minimising the negative effect in the charge distribution term by only having one $u$ quark. Together, these properties generate a large magnetic polarisability for the $\Xi^{0}$ and this is seen in our lattice QCD results.
The negatively charged baryons are very interesting. We observe that they have very small magnetic polarisability values as predicted by the quark model. However with the extrapolation incorporating divergent chiral physics and including both electro-quenching corrections and finite-volume corrections, both baryons are predicted to have positive values in the infinite-volume world of full QCD. Carefully examining the results in Tab. 7, we see there is a clear prediction for the $\Sigma^{-}$ to have a positive non-vanishing magnetic polarisability. On the other hand, the positive polarisability for the $\Xi^{-}$ is only a $1\,\sigma$ effect.
Figure 11: The valence-valence finite-volume (FV), electro-quenching (EQ), and total (FV+EQ) corrections for each outer octet baryon as a function of pion mass. These corrections are added to the original results from the lattice.
Also of interest is the very small difference between the values of the proton and neutron. Based on our discussion of the quark model in Sec. II, we would expect the neutron to have a larger magnetic polarisability. The additional up quark in the proton acts to increase the magnitude of the negative charge distribution term. Otherwise, the transition term is identical for the proton and neutron. We will examine this difference more carefully in Sec. X.
We now turn our attention to the subtle electro-quenching and finite-volume corrections for the baryons. The details of these corrections are shown in Fig. 11. With the exception of the valence-valence volume correction for the $\Sigma^{-}\,$, these corrections act to increase the magnetic polarisability. The $\Sigma^{-}\,$is unique and is discussed further below.
In all cases, the dominant correction at heavy pion masses is the electro-quenching correction. However, this correction is still small. As the pion mass approaches the physical value, both volume and electro-quenching corrections increase in magnitude, but the volume corrections increase much more quickly, resulting in corrections of approximately equal magnitude at the lightest pion mass considered in this work.
The $\Sigma^{-}\,$is interesting as it has large electro-quenching corrections. In the valence-valence sector, the sign of the chiral curvature is negative, such that the finite-volume correction of the valence-valence sector acts to decrease the magnetic polarisability. This is shown in the upper plot of Fig. 12. Upon implementing the electro-quenching correction, the sign of the chiral curvature becomes positive as illustrated in the lower plot of Fig. 12.
This change in sign underscores the importance of the order in which the corrections are implemented. Finite-volume corrections in the valence-valence sector are implemented first as this is the frame in which the finite-volume lattice results are obtained. Once volume corrected the lattice results can be fit to determine the coefficients of the residual series. Provided one uses a physical regulator parameter, the residual series coefficients are insensitive to sea-quark-loop contributions and one can model the electro-quenching corrections by changing the coefficients of the loop-integral coefficients. With the full QCD result in hand, one can plot the finite-volume- and electro-quenched-corrected points and draw the infinite-volume extrapolation curve. From here one can also explore other finite volumes as done in Fig. 10 and Fig. 12.
Figure 12: The chiral extrapolation of the finite-volume corrected valence-valence lattice QCD data (top) compared to the finite-volume and electro-quenching corrected extrapolation (bottom) for the $\Sigma^{-}\,$. The magnitude of the electro-quenching correction results in a sign change for the nonanalytic chiral curvature, a behaviour not observed for any other baryon.
IX Comparison to other predictions
The Particle Data Group Workman and Others (2022) quotes values for the magnetic polarisability of the proton and neutron. The values
$\beta_{p}^{\rm PDG}=2.5(4)\times 10^{-4}\,\text{fm}^{3},$ ((92))
$\beta_{n}^{\rm PDG}=3.7(12)\times 10^{-4}\,\text{fm}^{3},$ ((93))
are aggregated from a number of Compton scattering experiments. The quark model prediction that the neutron should have larger polarisability than the neutron is observed in the central values, though the large uncertainty of the neutron measurement precludes a definitive statement. Our final lattice QCD values reported in Tab. 7 compare favourably with the PDG values. Figure 13 shows the our chiral extrapolation in comparison to the PDG values and individual experiments for the proton and neutron.
Figure 13: The chiral extrapolation of volume- and electro-quenched-corrected lattice QCD results is compared with the PDG values Workman and Others (2022) and individual experiments Pasquini et al. (2019); MacGibbon et al. (1995); Blanpied et al. (2001); McGovern et al. (2013); de León et al. (2001); Beane et al. (2003); Kossert et al. (2002, 2003); Myers et al. (2014); Griesshammer et al. (2012) for the proton and neutron.
(a) $\Sigma^{+}\,$.
(b) $\Sigma^{-}\,$.
(c) $\Xi^{0}\,$.
(d) $\Xi^{-}\,$.
Figure 14: The chiral extrapolation of volume- and electro-quenched-corrected lattice QCD results are compared with other phenomenological estimates Aleksejevs and Barkanova (2011); Gobbi et al. (1996); Tanushi et al. (2001); Scoccola et al. (1996); Deshmukh and Tiburzi (2018) for the hyperons.
It is also interesting to place our results in the context of previous lattice QCD studies. Here we discuss the results of three works Lee et al. (2006); Chang et al. (2015); Bignell et al. (2020). We first note that Refs. Lee et al. (2006) and Chang et al. (2015) were carried out in the quenched approximation. In addition, Refs. Lee et al. (2006) and Chang et al. (2015) did not use a background-field corrected clover action to remove the additive mass renormalisation associated with the Wilson term when a background field is applied Bignell et al. (2019).
In the early work of Ref. Lee et al. (2006) the magnetic polarisabilities of all octet baryons were determined at a range of relatively heavy pion masses. In that work, the Landau term was not considered. However, we find the energy shift associated with the Landau term to be of similar magnitude to the total energy shift and therefore we do not compare with their results for charged baryons. Accounting for their different definition of the magnetic polarisability, we divide their results by $4\pi$ and compare the $\Xi^{0}\,$and neutron magnetic polarisabilities. We find our results for $\Xi^{0}\,$to be much larger than that reported in Ref. Lee et al. (2006), approaching a factor of 3. Our results for the neutron at the lightest quark mass considered in Ref. Lee et al. (2006) are 60% larger.
The observed discrepancy may be associated with the boundary conditions explored in Ref. Lee et al. (2006). While a fixed spatial boundary condition avoids the uniform magnetic field quantisation condition of Eq. (45), it is difficult to avoid the spatial boundary in the lattice QCD simulations introducing new systematic errors.
The lattice simulation of Ref. Chang et al. (2015) determined the polarisability of the proton and neutron at $m_{\pi}\sim 806\,$MeV. Again, the neutron results presented there are small at approximately half that presented here. In the case of the proton, they do attempt to fit the Landau term, using their smaller field strength correlation functions to identify which Landau level the particle rests in. They fit for the $n$th Landau level, allowing $n$ to take a positive real value. As $n$ should be an integer, they round the resulting value. However, due to the magnitude of the Landau term relative to the energy shift, their approach introduces a large uncertainty into the energy shift. As such, we do not consider their proton results further. Perhaps it is worthy to note, our use of Landau mode projection at the baryon sink removes all uncertainty surrounding the Landau levels.
Finally, Bignell et al. Bignell et al. (2020) determined the proton and neutron magnetic polarisabilities on the PACS-CS ensembles using similar methods which we have extended herein. The most important extensions include the single-state isolation analysis in all correlators, the consideration of several fit windows through weighted averaging, and much higher statistics. Our neutron results have a larger slope with respect to $m_{\pi}^{2}$ but extrapolate to the same value at the physical point. Increased statistics and associated methods have resolved smaller polarisability values for the proton at the two lightest quark masses considered, increasing the slope with respect to $m_{\pi}^{2}$ and producing smaller proton magnetic polarisabilities, bringing the previous value of $2.79(22)(16)\times 10^{-4}\,$fm3 to $2.12(17)(16)\times 10^{-4}\,$fm3, just outside of $1\sigma$ agreement.
Looking more broadly, there are several phenomenological studies of some of the hyperons. In the absence of experimental measurements, we compare our results with theoretical models in Fig. 14. It is here that one can observe the value of our lattice QCD simulation results for the hyperons. Our uncertainties are very small on the scale of variation in model predictions.
X Proton-Neutron Magnetic Polarisability Difference
The difference between the magnetic polarisability of the proton and neutron can provide a test of Reggeon dominance Gasser et al. (2015, 2020). Under the assumption of Reggeon dominance, chiral perturbation theory and Baldin sum rules can predict the difference of magnetic polarisabilities.
Figure 15: The difference between the proton and neutron magnetic polarisabilities. This work’s finite-volume and electro-quenched corrected lattice QCD results (FV & EQ Corr.) are extrapolated to the infinite volume (Inf. Vol.) physical point through chiral extrapolation. These results are compared to the experimental results of the PDG Workman and Others (2022), a Reggeon dominance prediction Gasser et al. (2020) and a previous lattice QCD calculation Bignell et al. (2020) which are horizontally offset at the physical pion mass for clarity.
We calculate the difference between the magnetic polarisability of the proton and neutron by construction of a correlation function ratio analogous to Eq. VI, which provides direct access to the polarisability difference
$\delta E_{\beta_{p}-\beta_{n}}$ $(B,t)=\frac{1}{2}\frac{1}{\delta t}\lim_{t\rightarrow\infty}\log\Bigl(\frac{R_{p}(B,t)}{R_{p}(B,t+\delta t)}\frac{R_{n}(B,t+\delta t)}{R_{n}(B,t)}\Bigr{missing}),$
[…]
$\delta E_{\beta_{p}-\beta_{n}}$ is fit using the techniques discussed in Sec. VI. We then extrapolate to the physical regime using the formalism discussed in Sec. VIII. When taking the polarisability difference, the $u-d$ symmetry in the leading loop-integral coefficients of the chiral expansion in full QCD causes the chiral contributions to cancel. As such, the extrapolation becomes a simple linear extrapolation. The resulting value at the physical point is
$\beta_{p}-\beta_{n}=0.09(11)\times 10^{-4}\,\textrm{fm}^{3}\,,$ ((95))
a much better estimate for the difference between the two polarisabilities. Here the statistical uncertainty is given in parentheses and the systematic uncertainty is negligible. Fig. 15 shows the extrapolation to the physical regime and includes the PDG value Workman and Others (2022), a result derived using Reggeon dominance Gasser et al. (2020) and a previous lattice QCD calculation Bignell et al. (2020).
The key observation is that our result is now in $1\sigma$ agreement with both experiment and the Reggeon dominance prediction, resolving a discrepancy observed in Ref. Bignell et al. (2020). Moreover, our results provide a very precise prediction for $\beta_{p}-\beta_{n}$.
XI Conclusion
A generalised expression for the magnetic polarisability was derived in a simple constituent quark model. Once corrected, this simple model provides accurate predictions for the octet-baryon magnetic polarisabilities examined herein. It also provides the ability to identify key characteristics of the magnetic polarisability and deep insight into the physics that drives the observed patterns. In particular, we identified the importance of opposite-charge quark flavours in generating large magnetic polarisability values. Here the importance of the up quark is manifest. We also focused on the octet-decuplet mass splitting and associated hyperfine interactions that govern the magnitude of the magnetic polarisability transition term.
Turning our attention to lattice simulation techniques, we investigated the behaviour of the underlying correlation functions associated with the correlation function ratio required to extract the magnetic polarisability in the background field method. We saw that in many cases, the correlation function ratio exhibits plateau-like behaviour in spite of the underlying correlation functions not yet having reached single-state isolation. We developed new methods to address this, ensuring the suppression of excited state contamination, and ensuring the uncertainties in our fitted mass shifts were not underestimated.
We also implemented a weighted averaging method to systematically extract the magnetic polarisability from the associated energy shifts obtained from candidate fit windows. This is especially important in cases where different fit windows introduce variation in the fitted values of the polarisability energy shifts. These improved analysis techniques allow the extraction of magnetic polarisability values for the outer octet baryons at a variety of pion masses with unprecedented precision.
The systematics of the lattice QCD calculation were then addressed. Drawing on chiral effective field theory and accounting for both electro-quenching and finite volume effects we extrapolated our simulation results to the physical point. This process produces results that compare favourably with the experimental values for the proton and neutron. Excellent agreement with the precise experimental value for the proton is observed and our prediction for the neutron is much more precise.
Comparison of phenomenological values for the hyperons indicate our results are very precise on the scale of variation in QCD models. Finally a new precise calculation of the difference in the proton and neutron magnetic polarisabilities has been presented.
We have revealed complex dynamics underpinning the magnetic polarisabilities of octet baryons. It would be interesting to examine these dynamics at a more microscopic level, where the quark mass dependence and environment dependence of individual quark sector contributions to the baryon polarisabilities are exposed. The techniques of the current presentation are flexible enough to admit such an analysis and we anticipate reporting on this in the near future.
Another challenge facing the community is gaining access to the light quark-mass regime. Using the techniques presented here, we are unable to resolve a signal of statistical interest. There are several issues that can lead to large uncontrolled statistical fluctuations and it will be important to examine this further to determine the cause of the fluctuations and perhaps develop new techniques to enable an exploration of the chiral regime.
Finally, the $uds$ members of the baryon octet, the $\Lambda$ and $\Sigma^{0}$ remain to be examined. As highlighted in the introduction, a perturbative calculation Wilcox and Lee (2021); Lee et al. (2023) can offer some important advantages in obtaining a clearer understanding of the magnetic polarisabilities of these baryons.
Acknowledgements.
We thank the PACS-CS Collaboration for making their $2+1$ flavour configurations available and the ongoing support of the International Lattice Data Grid (ILDG). Baryon correlation functions were constructed using the COLA software library, developed at the University of Adelaide Kamleh (2023). WK was supported by the Pawsey Supercomputing Centre through the Pawsey Centre for Extreme Scale Readiness (PaCER) program. RB is grateful for support via STFC grant ST/T000813/1 and acknowledges support from a Science Foundation Ireland Frontiers for the Future Project award with grant number SFI-21/FFP-P/10186 (data not available). This work was supported with supercomputing resources provided by the Phoenix HPC service at the University of Adelaide. This research was undertaken with the assistance of resources from the National Computational Infrastructure (NCI), which is supported by the Australian Government. This research is supported by Australian Research Council through Grants No. DP190102215 and No. DP210103706.
Appendix A Chiral coefficients
Table 8: Chiral coefficients $\chi^{2}_{MB}$ required to calculate quark-flow connected and disconnected coefficients for the leading chiral contribution to the magnetic polarisability. These coefficients address outer octet baryon transitions to an intermediate octet-baryon (column) octet-meson (row) states. Note that $\pi^{0}$ intermediate states do not contribute to the analysis. In full QCD the $\pi^{0}$ is electrically neutral and in partially quenched QCD, its composition of $\bar{u}u$ and $\bar{d}d$ matched flavours means it is not relevant to the disconnected flow calculations.
$p$ $n$ $\Xi^{0}$ $\Xi^{-}$ $\Sigma^{+}$ $\Sigma^{-}$ $\Sigma^{0}$ $\Lambda$
${\pi^{+}}$ $2\left(D+F\right)^{2}$ $2\left(D-F\right)^{2}$ $4F^{2}$ $\frac{4}{3}D^{2}$
$\pi^{-}$ $2\left(D+F\right)^{2}$ $2\left(D-F\right)^{2}$ $4F^{2}$ $\frac{4}{3}D^{2}$
${K^{+}}$ $2\left(D+F\right)^{2}$ $2\left(D-F\right)^{2}$ $\left(D-F\right)^{2}$ $\frac{1}{3}\left(D+3F\right)^{2}$
$K^{0}$ $2\left(D+F\right)^{2}$ $2\left(D-F\right)^{2}$ $\left(D-F\right)^{2}$ $\frac{1}{3}\left(D+3F\right)^{2}$
$\overline{K}^{0}$ $2\left(D-F\right)^{2}$ $2\left(D+F\right)^{2}$ $\left(D+F\right)^{2}$ $\frac{1}{3}\left(D-3F\right)^{2}$
$K^{-}$ $2\left(D-F\right)^{2}$ $2\left(D+F\right)^{2}$ $\left(D+F\right)^{2}$ $\frac{1}{3}\left(D-3F\right)^{2}$
Table 9: Chiral coefficients $\chi^{2}_{MB}$ required to calculate quark-flow connected and disconnected coefficients for the leading chiral contribution to the magnetic polarisability. These coefficients address outer octet baryon transitions to an intermediate decuplet-baryon (column) octet-meson (row) states. Again $\pi^{0}$ intermediate states do not contribute.
$\Delta^{++}$ $\Delta^{+}$ $\Delta^{0}$ $\Delta^{-}$ $\Sigma^{*+}$ $\Sigma^{*0}$ $\Sigma^{*-}$ $\Xi^{*0}$ $\Xi^{*-}$ $\Omega^{-}$
${\pi^{+}}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{4}{3}\mathcal{C}^{2}$ $\frac{2}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$
$\pi^{-}$ $\frac{4}{3}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{2}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$
${K^{+}}$ $\frac{2}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{4}{3}\mathcal{C}^{2}$
$K^{0}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{2}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{4}{3}\mathcal{C}^{2}$
$\overline{K}^{0}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{8}{9}\mathcal{C}^{2}$ $\frac{4}{3}\mathcal{C}^{2}$ $\frac{2}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$
$K^{-}$ $\frac{4}{3}\mathcal{C}^{2}$ $\frac{8}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{4}{9}\mathcal{C}^{2}$ $\frac{2}{9}\mathcal{C}^{2}$
Table 10: Chiral coefficients for the leading-order loop integral contributions for the neutron.
Process Total Valence-sea Sea-sea
$n\rightarrow N\pi$
$n\rightarrow n\pi^{0}$ 00 $2q_{d}q_{\bar{d}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})+2q_{u}q_{\bar{u}}\,\chi^{2}_{K^{+}\Sigma^{-}}$ $q^{2}_{\bar{d}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})+q^{2}_{\bar{u}}\,\chi^{2}_{K^{+}\Sigma^{-}}$
$n\rightarrow p\pi^{-}$ $\chi^{2}_{\pi^{-}p}$ $2q_{d}q_{\bar{u}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$ $q^{2}_{\bar{u}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$
$n\rightarrow n^{-}\pi^{+}$ 00 $2q_{u}q_{\bar{d}}\,\chi^{2}_{K^{+}\Sigma^{-}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{+}\Sigma^{-}}$
$n\rightarrow\Sigma K$
$n\rightarrow(\Sigma^{0},\,\Lambda)K^{0}$ 00 $2q_{d}q_{\bar{s}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$ $q^{2}_{\bar{s}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$
$n\rightarrow\Sigma^{-}K^{+}$ $\chi^{2}_{K^{+}\Sigma^{-}}$ $2q_{u}q_{\bar{s}}\,\chi^{2}_{K^{+}\Sigma^{-}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{K^{+}\Sigma^{-}}$
$n\rightarrow\Delta\pi$
$n\rightarrow\Delta^{0}\pi^{0}$ 00 $2q_{d}q_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{*0}}+2q_{u}q_{\bar{u}}\,\chi^{2}_{K^{+}\Sigma^{*-}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{*0}}+q^{2}_{\bar{u}}\,\chi^{2}_{K^{+}\Sigma^{*-}}$
$n\rightarrow\Delta^{+}\pi^{-}$ $\chi^{2}_{\pi^{-}\Delta^{+}}$ $2q_{d}q_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$
$n\rightarrow\Delta^{-}\pi^{+}$ $\chi^{2}_{\pi^{+}\Delta^{-}}$ $2q_{u}q_{\bar{d}}\,\chi^{2}_{K^{+}\Sigma^{*-}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{+}\Sigma^{*-}}$
$n\rightarrow\Sigma^{*}K$
$n\rightarrow\Sigma^{*0}K^{0}$ 00 $2q_{d}q_{\bar{s}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$
$n\rightarrow\Sigma^{*-}K^{+}$ $\chi^{2}_{K^{+}\Sigma^{*-}}$ $2q_{u}q_{\bar{s}}\,\chi^{2}_{K^{+}\Sigma^{*-}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{K^{+}\Sigma^{*-}}$
Table 11: Chiral coefficients for the leading-order loop integral contributions for the proton.
Process Total Valence-sea Sea-sea
$p\rightarrow N\pi$
$p\rightarrow p\pi^{0}$ 00 $2q_{u}q_{\bar{u}}(\chi^{2}_{K^{+}\Sigma^{0}}+\chi^{2}_{K^{+}\Lambda})+2q_{d}q_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{+}}$ $q^{2}_{\bar{u}}(\chi^{2}_{K^{+}\Sigma^{0}}+\chi^{2}_{K^{+}\Lambda})+q^{2}_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{+}}$
$p\rightarrow n\pi^{+}$ $\chi^{2}_{\pi^{+}n}$ $2q_{u}q_{\bar{d}}(\chi^{2}_{K^{+}\Sigma^{0}}+\chi^{2}_{K^{+}\Lambda})$ $q^{2}_{\bar{d}}(\chi^{2}_{K^{+}\Sigma^{0}}+\chi^{2}_{K^{+}\Lambda})$
$p\rightarrow p^{++}\pi^{-}$ 00 $2q_{d}q_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{+}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{+}}$
$p\rightarrow\Sigma K$
$p\rightarrow(\Sigma^{0},\,\Lambda)K^{+}$ $\chi^{2}_{K^{+}\Sigma^{0}}+\chi^{2}_{K^{+}\Lambda}$ $2q_{u}q_{\bar{s}}(\chi^{2}_{K^{+}\Sigma^{0}}+\chi^{2}_{K^{+}\Lambda})$ $q^{2}_{\bar{s}}(\chi^{2}_{K^{+}\Sigma^{0}}+\chi^{2}_{K^{+}\Lambda})$
$p\rightarrow\Sigma^{+}K^{0}$ 00 $2q_{d}q_{\bar{s}}\,\chi^{2}_{K^{0}\Sigma^{+}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{K^{0}\Sigma^{+}}$
$p\rightarrow\Delta\pi$
$p\rightarrow\Delta^{+}\pi^{0}$ 00 $2q_{u}q_{\bar{u}}\,\chi^{2}_{K^{+}\Sigma^{*0}}+2q_{d}q_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{*+}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{+}\Sigma^{*0}}+q^{2}_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{*+}}$
$p\rightarrow\Delta^{0}\pi^{+}$ $\chi^{2}_{\pi^{+}\Delta^{0}}$ $2q_{u}q_{\bar{d}}\,\chi^{2}_{K^{+}\Sigma^{*0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{+}\Sigma^{*0}}$
$p\rightarrow\Delta^{++}\pi^{-}$ $\chi^{2}_{\pi^{-}\Delta^{++}}$ $2q_{d}q_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{*+}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{*+}}$
$p\rightarrow\Sigma^{*}K$
$p\rightarrow\Sigma^{*0}K^{+}$ $\chi^{2}_{K^{+}\Sigma^{*0}}$ $2q_{u}q_{\bar{s}}\,\chi^{2}_{K^{+}\Sigma^{*0}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{K^{+}\Sigma^{*0}}$
$p\rightarrow\Sigma^{*+}K^{0}$ 00 $2q_{d}q_{\bar{s}}\,\chi^{2}_{K^{0}\Sigma^{*+}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{K^{0}\Sigma^{*+}}$
Table 12: Chiral coefficients for the leading-order loop integral contributions for the $\Sigma^{+}$.
Process Total Valence-sea Sea-sea
$\Sigma^{+}\rightarrow\Sigma\pi$
$\Sigma^{+}\rightarrow\Sigma^{+}\pi^{0}$ 00 $2q_{u}q_{\bar{u}}(\chi^{2}_{\pi^{+}\Sigma^{0}}+\chi^{2}_{\pi^{+}\Lambda})$ $q^{2}_{\bar{u}}(\chi^{2}_{\pi^{+}\Sigma^{0}}+\chi^{2}_{\pi^{+}\Lambda})$
$\Sigma^{+}\rightarrow(\Sigma^{0},\,\Lambda)\pi^{+}$ $\chi^{2}_{\pi^{+}\Sigma^{0}}+\chi^{2}_{\pi^{+}\Lambda}$ $2q_{u}q_{\bar{d}}(\chi^{2}_{\pi^{+}\Sigma^{0}}+\chi^{2}_{\pi^{+}\Lambda})$ $q^{2}_{\bar{d}}(\chi^{2}_{\pi^{+}\Sigma^{0}}+\chi^{2}_{\pi^{+}\Lambda})$
$\Sigma^{+}\rightarrow NK$
$\Sigma^{+}\rightarrow p^{++}K^{-}$ 00 $2q_{s}q_{\bar{u}}\,\chi^{2}_{K^{0}p}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{0}p}$
$\Sigma^{+}\rightarrow pK^{0}$ 00 $2q_{s}q_{\bar{d}}\,\chi^{2}_{K^{0}p}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{0}p}$
$\Sigma^{+}\rightarrow\Xi K$
$\Sigma^{+}\rightarrow\Xi^{0}K^{+}$ $\chi^{2}_{K^{+}\Xi^{0}}$ $2q_{u}q_{\bar{s}}(\chi^{2}_{\pi^{+}\Sigma^{0}}+\chi^{2}_{\pi^{+}\Lambda})$ $q^{2}_{\bar{s}}(\chi^{2}_{\pi^{+}\Sigma^{0}}+\chi^{2}_{\pi^{+}\Lambda})$
$\Sigma^{+}\rightarrow\Sigma^{*}\pi$
$\Sigma^{+}\rightarrow\Sigma^{*+}\pi^{0}$ 00 $2q_{u}q_{\bar{u}}\,\chi^{2}_{\pi^{+}\Sigma^{*0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{\pi^{+}\Sigma^{*0}}$
$\Sigma^{+}\rightarrow\Sigma^{*0}\pi^{+}$ $\chi^{2}_{\pi^{+}\Sigma^{*0}}$ $2q_{u}q_{\bar{d}}\,\chi^{2}_{\pi^{+}\Sigma^{*0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{\pi^{+}\Sigma^{*0}}$
$\Sigma^{+}\rightarrow\Delta K$
$\Sigma^{+}\rightarrow\Delta^{++}K^{-}$ $\chi^{2}_{K^{-}\Delta^{++}}$ $2q_{s}q_{\bar{u}}\,\chi^{2}_{K^{0}\Delta^{+}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{0}\Delta^{+}}$
$\Sigma^{+}\rightarrow\Delta^{+}K^{0}$ 00 $2q_{s}q_{\bar{d}}\,\chi^{2}_{K^{0}\Delta^{+}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{0}\Delta^{+}}$
$\Sigma^{+}\rightarrow\Xi^{*}K$
$\Sigma^{+}\rightarrow\Xi^{*0}K^{+}$ $\chi^{2}_{K^{+}\Xi^{*0}}$ $2q_{u}q_{\bar{s}}\,\chi^{2}_{\pi^{+}\Sigma^{*0}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{\pi^{+}\Sigma^{*0}}$
Table 13: Chiral coefficients for the leading-order loop integral contributions for the $\Xi^{0}$.
Process Total Valence-sea Sea-sea
$\Xi^{0}\rightarrow\Xi\pi$
$\Xi^{0}\rightarrow\Xi^{0}\pi^{0}$ 00 $2q_{u}q_{\bar{u}}\,\chi^{2}_{\pi^{+}\Xi^{-}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{\pi^{+}\Xi^{-}}$
$\Xi^{0}\rightarrow\Xi^{-}\pi^{+}$ $\chi^{2}_{\pi^{+}\Xi^{-}}$ $2q_{u}q_{\bar{d}}\,\chi^{2}_{\pi^{+}\Xi^{-}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{\pi^{+}\Xi^{-}}$
$\Xi^{0}\rightarrow\Sigma K$
$\Xi^{0}\rightarrow\Sigma^{+}K^{-}$ $\chi^{2}_{K^{-}\Sigma^{+}}$ $2q_{s}q_{\bar{u}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$ $q^{2}_{\bar{u}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$
$\Xi^{0}\rightarrow(\Sigma^{0},\,\Lambda)K^{0}$ 00 $2q_{s}q_{\bar{d}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$ $q^{2}_{\bar{d}}(\chi^{2}_{K^{0}\Sigma^{0}}+\chi^{2}_{K^{0}\Lambda})$
$\Xi^{0}\rightarrow\Xi K$
$\Xi^{0}\rightarrow\Xi^{-}_{3s}K^{+}$ 00 $2q_{u}q_{\bar{s}}\,\chi^{2}_{\pi^{+}\Xi^{-}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{\pi^{+}\Xi^{-}}$
$\Xi^{0}\rightarrow\Xi^{*}\pi$
$\Xi^{0}\rightarrow\Xi^{*0}\pi^{0}$ 00 $2q_{u}q_{\bar{u}}\,\chi^{2}_{\pi^{+}\Xi^{*-}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{\pi^{+}\Xi^{*-}}$
$\Xi^{0}\rightarrow\Xi^{*-}\pi^{+}$ $\chi^{2}_{\pi^{+}\Xi^{*-}}$ $2q_{u}q_{\bar{d}}\,\chi^{2}_{\pi^{+}\Xi^{*-}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{\pi^{+}\Xi^{*-}}$
$\Xi^{0}\rightarrow\Sigma^{*}K$
$\Xi^{0}\rightarrow\Sigma^{*+}K^{-}$ $\chi^{2}_{K^{-}\Sigma^{*+}}$ $2q_{s}q_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$
$\Xi^{0}\rightarrow\Sigma^{*0}K^{0}$ 00 $2q_{s}q_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{0}\Sigma^{*0}}$
$\Xi^{0}\rightarrow\Omega K$
$\Xi^{0}\rightarrow\Omega^{-}K^{+}$ $\chi^{2}_{K^{+}\Omega^{-}}$ $2q_{u}q_{\bar{s}}\,\chi^{2}_{\pi^{+}\Xi^{*-}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{\pi^{+}\Xi^{*-}}$
Table 14: Chiral coefficients for the leading-order loop integral contributions for the $\Sigma^{-}$.
Process Total Valence-sea Sea-sea
$\Sigma^{-}\rightarrow\Sigma\pi$
$\Sigma^{-}\rightarrow\Sigma^{-}\pi^{0}$ 00 $2q_{d}q_{\bar{d}}(\chi^{2}_{\pi^{-}\Sigma^{0}}+\chi^{2}_{\pi^{-}\Lambda})$ $q^{2}_{\bar{d}}(\chi^{2}_{\pi^{-}\Sigma^{0}}+\chi^{2}_{\pi^{-}\Lambda})$
$\Sigma^{-}\rightarrow(\Sigma^{0},\,\Lambda)\pi^{-}$ $\chi^{2}_{\pi^{-}\Sigma^{0}}+\chi^{2}_{\pi^{-}\Lambda}$ $2q_{d}q_{\bar{u}}(\chi^{2}_{\pi^{-}\Sigma^{0}}+\chi^{2}_{\pi^{-}\Lambda})$ $q^{2}_{\bar{u}}(\chi^{2}_{\pi^{-}\Sigma^{0}}+\chi^{2}_{\pi^{-}\Lambda})$
$\Sigma^{-}\rightarrow NK$
$\Sigma^{-}\rightarrow nK^{-}$ $\chi^{2}_{K^{-}n}$ $2q_{s}q_{\bar{u}}\,\chi^{2}_{K^{-}n}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{-}n}$
$\Sigma^{-}\rightarrow n^{-}K^{0}$ 00 $2q_{s}q_{\bar{d}}\,\chi^{2}_{K^{-}n}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{-}n}$
$\Sigma^{-}\rightarrow\Xi K$
$\Sigma^{-}\rightarrow\Xi^{-}K^{0}$ 00 $2q_{d}q_{\bar{s}}(\chi^{2}_{\pi^{-}\Sigma^{0}}+\chi^{2}_{\pi^{-}\Lambda})$ $q^{2}_{\bar{s}}(\chi^{2}_{\pi^{-}\Sigma^{0}}+\chi^{2}_{\pi^{-}\Lambda})$
$\Sigma^{-}\rightarrow\Sigma^{*}\pi$
$\Sigma^{-}\rightarrow\Sigma^{*-}\pi^{0}$ 00 $2q_{d}q_{\bar{d}}\,\chi^{2}_{\pi^{-}\Sigma^{*0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{\pi^{-}\Sigma^{*0}}$
$\Sigma^{-}\rightarrow\Sigma^{*0}\pi^{-}$ $\chi^{2}_{\pi^{-}\Sigma^{*0}}$ $2q_{d}q_{\bar{u}}\,\chi^{2}_{\pi^{-}\Sigma^{*0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{\pi^{-}\Sigma^{*0}}$
$\Sigma^{-}\rightarrow\Delta K$
$\Sigma^{-}\rightarrow\Delta^{0}K^{-}$ $\chi^{2}_{K^{-}\Delta^{0}}$ $2q_{s}q_{\bar{u}}\,\chi^{2}_{K^{-}\Delta^{0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{-}\Delta^{0}}$
$\Sigma^{-}\rightarrow\Delta^{-}K^{0}$ 00 $2q_{s}q_{\bar{d}}\,\chi^{2}_{K^{-}\Delta^{0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{-}\Delta^{0}}$
$\Sigma^{-}\rightarrow\Xi^{*}K$
$\Sigma^{-}\rightarrow\Xi^{*-}K^{0}$ 00 $2q_{d}q_{\bar{s}}\,\chi^{2}_{\pi^{-}\Sigma^{*0}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{\pi^{-}\Sigma^{*0}}$
Table 15: Chiral coefficients for the leading-order loop integral contributions for the $\Xi^{-}$.
Process Total Valence-sea Sea-sea
$\Xi^{-}\rightarrow\Xi\pi$
$\Xi^{-}\rightarrow\Xi^{-}\pi^{0}$ 00 $2q_{d}q_{\bar{d}}\,\chi^{2}_{\pi^{-}\Xi^{0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{\pi^{-}\Xi^{0}}$
$\Xi^{-}\rightarrow\Xi^{0}\pi^{-}$ $\chi^{2}_{\pi^{-}\Xi^{0}}$ $2q_{d}q_{\bar{u}}\,\chi^{2}_{\pi^{-}\Xi^{0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{\pi^{-}\Xi^{0}}$
$\Xi^{-}\rightarrow\Sigma K$
$\Xi^{-}\rightarrow\Sigma^{-}K^{0}$ 00 $2q_{s}q_{\bar{d}}(\chi^{2}_{K^{-}\Sigma^{0}}+\chi^{2}_{K^{-}\Lambda})$ $q^{2}_{\bar{d}}(\chi^{2}_{K^{-}\Sigma^{0}}+\chi^{2}_{K^{-}\Lambda})$
$\Xi^{-}\rightarrow(\Sigma^{0},\,\Lambda)K^{-}$ $\chi^{2}_{K^{-}\Sigma^{0}}+\chi^{2}_{K^{-}\Lambda}$ $2q_{s}q_{\bar{u}}(\chi^{2}_{K^{-}\Sigma^{0}}+\chi^{2}_{K^{-}\Lambda})$ $q^{2}_{\bar{u}}(\chi^{2}_{K^{-}\Sigma^{0}}+\chi^{2}_{K^{-}\Lambda})$
$\Xi^{-}\rightarrow\Xi K$
$\Xi^{-}\rightarrow\Xi^{-}_{3s}K^{0}$ 00 $2q_{d}q_{\bar{s}}\,\chi^{2}_{\pi^{-}\Xi^{0}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{\pi^{-}\Xi^{0}}$
$\Xi^{-}\rightarrow\Xi^{*}\pi$
$\Xi^{-}\rightarrow\Xi^{*-}\pi^{0}$ 00 $2q_{d}q_{\bar{d}}\,\chi^{2}_{\pi^{-}\Xi^{*0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{\pi^{-}\Xi^{*0}}$
$\Xi^{-}\rightarrow\Xi^{*0}\pi^{-}$ $\chi^{2}_{\pi^{-}\Xi^{*0}}$ $2q_{d}q_{\bar{u}}\,\chi^{2}_{\pi^{-}\Xi^{*0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{\pi^{-}\Xi^{*0}}$
$\Xi^{-}\rightarrow\Sigma^{*}K$
$\Xi^{-}\rightarrow\Sigma^{*-}K^{0}$ 00 $2q_{s}q_{\bar{d}}\,\chi^{2}_{K^{-}\Sigma^{*0}}$ $q^{2}_{\bar{d}}\,\chi^{2}_{K^{-}\Sigma^{*0}}$
$\Xi^{-}\rightarrow\Sigma^{*0}K^{-}$ $\chi^{2}_{K^{-}\Sigma^{*0}}$ $2q_{s}q_{\bar{u}}\,\chi^{2}_{K^{-}\Sigma^{*0}}$ $q^{2}_{\bar{u}}\,\chi^{2}_{K^{-}\Sigma^{*0}}$
$\Xi^{-}\rightarrow\Omega K$
$\Xi^{-}\rightarrow\Omega^{-}K^{0}$ 00 $2q_{d}q_{\bar{s}}\,\chi^{2}_{\pi^{-}\Xi^{*0}}$ $q^{2}_{\bar{s}}\,\chi^{2}_{\pi^{-}\Xi^{*0}}$
As discussed in Sec. VIII, calculating a correction for electro-quenching requires identification of the various valence-valence, valence-sea and sea-sea contributions to the magnetic polarisability. These contributions may be written in terms of the SU(3) chiral coefficients. The coefficients $\chi^{2}_{MB}$ are derived by considering an appropriate chiral Lagrangian. The reader is directed to Ref. Shanahan et al. (2013) for a convenient catalogue of these coefficients. The coefficients are written in terms of the SU(6) values $D,F,\mathcal{C}$. We take $D+F=g_{A}=1.267$, $F=\frac{2}{3}D$, $\mathcal{C}=-1.52$ Bignell et al. (2020).
The chiral coefficients for standard SU(3) axial transitions Shanahan et al. (2013) are listed in Tab. 8 and Tab. 9. Note, our definition of the coefficients in the loop integrals differ from Ref. Shanahan et al. (2013) such that their octet coefficients have been squared and divided by 2 and their decuplet coefficients have been squared and multiplied by $4/3$.
The contributions for each possible process are derived in the manner described in Sec. VIII. All contributions are listed in Tables 10 through 15.